### v1

# Beyond Sociodemographic Prompting: Using Supervision to Align LLMs with Human Response Distributions

Abstract
The ability to accurately predict how different population groups would answer subjective questions would have great value. In this work, we show that use of relatively simple supervision can greatly improve language model alignment with diverse population groups, as measured over three datasets spanning various topics. Beyond evaluating average performance, we also report how alignment varies across specific groups. The simplicity and generality of our approach promotes easy adoption, while our broad findings provide useful guidance for when to use or not use our approach in practice. By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a useful benchmark to stimulate future research 111Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment.
1 Introduction
Human surveys elicit valuable public opinion on a wide range of topics, such as proposed legislation, candidates for office (i.e., election polls), marketing or outreach campaigns, commercial products and services, etc. Hayati et al. (2024); Santurkar et al. (2023). Subjective annotation tasks in NLP also seek diverse, human judgments Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). Given the frequency, breadth, and importance of obtaining opinion data from diverse populations, there has been great interest in developing effective ways to conduct such surveys accurately and efficiently.
Recent work has sought to simulate human responses using LLMs in many areas, such as in psychology, sociology, and economic studies Aher et al. (2023), NLP annotation Nasution and Onan (2024), and large-scale survey creation and testing Rothschild et al. (2024). More specifically, LLMs are being increasingly used to simulate social or demographic groups Hu and Collier (2024). This is commonly done with persona or sociodemographic (SD) prompting Hu and Collier (2024); Beck et al. (2024) by both technical and non-technical practitioners alike Beck et al. (2024).
Prior work on SD prompting has typically suffered from one or more key shortcomings. First, many studies have assumed a single majority answer for each sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025); Mukherjee et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. In addition, prompting for a distribution, rather than a representative response, could potentially make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2024); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024). Other work seeking to usefully replicate ground-truth answer distributions have been agnostic to underlying demographics, i.e., evaluating LLM distributions prompted with specific demographics with ground-truth distributions of multiple other demographics Beck et al. (2024).
In this work, we evaluate SD prompting in a distributional manner across three large survey datasets (both US and global), using a variety of methods to extract distributions from LLMs. We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs. We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups.
Concretely, we evaluate and explore distributional opinion alignment by studying the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data? Is this consistent across models, datasets, and distribution elicitation methods? (RQ 3) Can generated distributions be more easily aligned with some SD groups over others? What is the effect of calibration for individual groups? (RQ 4) Does calibration work with fewer supervised training examples? And (RQ 5) What are the effects of post-training (methods like SFT and DPO) on LLM-generated distributions?
Our experiments evaluate the alignment of LLM-generated distributions to human opinion, across three distribution elicitation methods, and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We find that while baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1), our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across settings (RQ2). Moreover, as few as 5 gold examples can be used to calibrate distributions (RQ4), though alignment to some demographics is degraded more than others (RQ3). In regard to LLM post-training (RQ5), the effect appears to be dataset dependent, though we find that calibrated alignment is less affected than uncalibrated distributions.
We conclude with concrete suggestions to practitioners seeking to predict diverse human responses (e.g., public opinion surveys) via LLMs, offering appropriate caution.
2 Related Work
Sociodemographic (SD) Prompting
To support safe deployment of models, it is important that modeling and evaluation practices account for social dimensions of diversity Santurkar et al. (2023) that shape our social identities, including sociodemographic factors such as age, gender, and ethnicity Hwang et al. (2023); Movva et al. (2024). SD prompting refers to incorporating demographic variables into input prompts to enhance model alignment with desired outputs Alipour et al. (2024); Joshi et al. (2025). As past work has shown that neither annotator agreement methodologies nor reward mechanisms often produce significant differences in the solution Yang et al. (2024), prompting methods may support more efficient processes. Some work has reported that SD prompting can be effective in improving model performance Beck et al. (2024), supporting SD-specific diversification of predictions without need for collecting sensitive annotator SD information Gupta et al. (2023).
Pitfalls of Sociodemographic Prompting
Prior studies evaluating sociodemographic prompting have reported mixed results in effectiveness in accurately simulating a social group Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024). While some works show that sociodemographic prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024).
Other work warns against such methods due to misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2024); Cheng et al. (2023b). Simply incorporating demographic features into prompts does not always enable LLMs to adopt the perspectives of specific demographic groups Sun et al. (2025). Furthermore, a vast body of work has established biases in language models that may be perpetuated with sociodemographic prompting Nadeem et al. (2020); Gallegos et al. (2024); Gupta et al. (2024). There is risk of "pigeonholing particular demographic groups into specific narratives" Cheng et al. (2023a).
Distribution Elicitation
Confidence elicitation is the process of estimating LLM’s confidence in their responses without model fine-tuning or accessing internal information Geng et al. (2024a). Recent research has explored methods for confidence elicitation which are more suitable for both closed-source commercial APIs as well as open-source models Lin et al. (2024). Similarly, distribution elicitation methods are techniques used to obtain probability distributions from experts about unknown quantities Falconer et al. (2022).
3 Approach
3.1 Datasets
We consider three human survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA), and the World Values Survey (WVS). Both OQA and WVS have been used to study sociodemographic prompting in prior work Santurkar et al. (2023); Durmus et al. (2024). These three datasets cover a diverse range of subjective topics, such as perceptions of science and public health, public opinion on gun control, data privacy, and various moral opinions and values. To reduce downstream LLM costs, we subset datasets to include all WGM ordinal questions and two questions per category from OQA and WVS. This totals 92 questions across WGM (14), OQA (38), and WVS (40).
Demographics for each dataset are shown in Table 11 (Appendix). Over all datasets, questions, and demographics, there are 4,500 human response distributions. For each question, model, and distribution elicitation method (Section 3.2), we predict the probability distribution of human responses over answer choices. This yields 220,500 generated SD-specific response distributions and 4,500 generated SD-agnostic response distributions.
3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To predict these distributions via LLM, we apply (with some modifications) methods inspired from confidence literature as follows. We investigate techniques that work most broadly, supporting use with both black-box and white-box LLMs. That said, Appendix A.3 reports a smaller scale study with log probabilities. Appendix B shows prompts used for all methods.
Verbalized. This is simply directly prompting the model to output a distribution in numbers Geng et al. (2024b); Tian et al. (2023). For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over the answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample times and average for variance reduction. If a model ever does not generate a distribution as requested, we simply discard that output. We also renormalize generated distributions that do not sum to 1.
Self-random. We prompt the model to output a single answer choice, sampling times Xiong et al. (2024), with a temperature of 0.7. We then create a distribution over the responses. For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice, such as 4 for a lot or 1 for not at all. Note that this method estimates the log probabilities for large Tian et al. (2023). In our experiments, we sample times.
Paraphrase. We also prompt the model to output a single answer choice but use different paraphrases of the prompt instead of sampling responses for the same prompt Xiong et al. (2024). We use paraphrases for each prompt.
3.3 Prompts
We use two categories of prompts, and for each category, we prompt with the different elicitation methods. See Appendix B for prompts used.
Base. Base prompts exclude any sociodemographic information, asking questions from the original dataset with formatting slightly changed for LLMs. See Appendix Table 12 for full prompts.
Sociodemographic (SD). These use a form similar to “Imagine you are {d}. Question: ”, where {d} is a demographic value. The exact prompt varies with demographic and elicitation method. See Appendix Table 13 for full prompts used.
3.4 Metric
We use the opinion alignment metric from Santurkar et al. (2023) to measure similarity between elicited distributions and ground truth distributions for all results. This metric makes use of 1 Wasserstein distance, or earth-mover’s distance. As noted by Santurkar et al. (2023), this metric takes into account the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges from 0 to 1, but in our results we show the value as a percentage.
3.5 Models
We evaluate 15 models ranging in openness (open-source, open-weight, black box), size, modality, and post-training. Model families include Claude, Llama, Mistral, OLMo-2, and Qwen. Our main results report on only the most powerful model in each family. Others are reported in Appendix A.1.
3.6 Calibration
We apply supervised regression to transform LLM-generated distributions. We split the LLM distributions into train, development, and test sets (60-20-20). Results in Section 6 show that far less supervision is typically needed in practice.
The input to regression is the LLM distribution and the output is the scaled distribution. We split the distributions by answer choice before normalizing across choices. More formally, for a regression model and LLM distribution parameterized by the values for each answer choice , we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice . I.e., we train on pairs of to learn transformed values on held-out test questions. We learn a regression model for each dataset-LLM-elicitation setting and do development set hyperparameter tuning and model selection to choose between linear regression (including lasso and ridge) and random forest regression.
| Model | Base prompt | Sociodemographic prompt | |||||||||||
| WGM | Claude-3.5-v2 | 66.2 | 85.2 | 59.5 | 85.1 | 89.3 | 87.7 | 65.4 | 84.4 | 61.4 | 84.2 | 89.0 | 89.3 |
| Llama-3.2-90B | 68.1 | 84.6 | 73.0 | 86.2 | 84.8 | 89.0 | 70.6 | 86.0 | 67.6 | 86.1 | 85.0 | 89.8 | |
| Mistral-large | 62.4 | 84.7 | 72.0 | 83.9 | 89.4 | 88.9 | 68.4 | 84.7 | 63.0 | 84.9 | 87.3 | 88.4 | |
| OLMo-2-7B-I | 59.6 | 81.5 | 67.7 | 80.4 | 59.8 | 85.1 | 62.3 | 82.5 | 64.5 | 82.7 | 69.8 | 84.6 | |
| Qwen-2.5-72B | 57.7 | 83.1 | 53.3 | 83.2 | 88.2 | 87.1 | 66.0 | 84.1 | 63.4 | 85.2 | 89.1 | 89.4 | |
| Average | 62.8 | 83.8 | 65.1 | 83.8 | 82.3 | 87.6 | 66.5 | 84.3 | 64.0 | 84.6 | 84.0 | 88.3 | |
| OQA | Claude-3.5-v2 | 70.5 | 88.8 | 72.5 | 90.4 | 91.7 | 91.7 | 76.1 | 89.9 | 73.3 | 89.6 | 91.9 | 91.6 |
| Llama-3.2-90B | 79.3 | 89.6 | 75.3 | 89.4 | 86.8 | 87.9 | 79.2 | 90.1 | 76.1 | 90.0 | 83.4 | 85.9 | |
| Mistral-large | 79.5 | 89.9 | 75.3 | 88.3 | 85.0 | 86.2 | 75.8 | 89.2 | 72.4 | 89.5 | 83.8 | 84.7 | |
| OLMo-2-7B-I | 72.6 | 89.0 | 72.3 | 88.4 | 65.4 | 79.9 | 72.8 | 88.4 | 70.5 | 88.6 | 68.5 | 81.5 | |
| Qwen-2.5-72B | 73.9 | 89.7 | 67.0 | 88.8 | 88.4 | 87.9 | 74.4 | 90.0 | 71.3 | 89.5 | 89.2 | 88.6 | |
| Average | 75.2 | 89.4 | 72.5 | 89.1 | 83.5 | 86.7 | 75.7 | 89.5 | 72.7 | 89.4 | 83.4 | 86.5 | |
| WVS | Claude-3.5-v2 | 46.5 | 80.3 | 51.0 | 80.3 | 75.2 | 80.3 | 61.0 | 80.4 | 56.8 | 80.4 | 75.6 | 81.7 |
| Llama-3.2-90B | 61.6 | 79.9 | 59.1 | 80.3 | 64.6 | 80.3 | 62.1 | 80.8 | 59.5 | 81.5 | 67.7 | 82.7 | |
| Mistral-large | 48.8 | 80.3 | 44.0 | 82.2 | 72.8 | 80.3 | 54.3 | 80.4 | 51.5 | 80.3 | 76.6 | 83.8 | |
| OLMo-2-7B-I | 75.3 | 80.3 | 74.9 | 80.3 | 86.6 | 86.5 | 58.3 | 79.2 | 60.0 | 77.2 | 86.0 | 89.8 | |
| Qwen-2.5-72B | 39.6 | 82.0 | 42.4 | 82.3 | 74.0 | 77.9 | 49.1 | 82.4 | 49.4 | 81.5 | 73.4 | 81.7 | |
| Average | 54.4 | 80.6 | 54.3 | 81.1 | 74.6 | 81.1 | 57.0 | 80.6 | 55.4 | 80.2 | 75.9 | 83.9 | |
| Average | Claude-3.5-v2 | 61.1 | 84.8 | 61.0 | 85.3 | 85.4 | 86.6 | 67.5 | 84.9 | 63.8 | 84.7 | 85.5 | 87.5 |
| Llama-3.2-90B | 69.7 | 84.7 | 69.1 | 85.3 | 78.7 | 85.7 | 70.6 | 85.6 | 67.7 | 85.9 | 78.7 | 86.1 | |
| Mistral-large | 63.6 | 85.0 | 63.8 | 84.8 | 82.4 | 85.1 | 66.2 | 84.8 | 62.3 | 84.9 | 82.6 | 85.6 | |
| OLMo-2-7B-I | 69.2 | 83.6 | 71.6 | 83.0 | 70.6 | 83.8 | 64.5 | 83.4 | 65.0 | 82.8 | 74.8 | 85.3 | |
| Qwen-2.5-72B | 57.1 | 84.9 | 54.2 | 84.8 | 83.5 | 84.3 | 63.2 | 85.5 | 61.4 | 85.4 | 83.9 | 86.6 | |
| Average | 64.1 | 84.6 | 64.0 | 84.6 | 80.1 | 85.1 | 66.4 | 84.8 | 64.0 | 84.7 | 81.1 | 86.2 | |
4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts, and calibrating LLM-generated distributions using some human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The full results for all models are shown in Appendix Table 5.
Findings (RQ1, RQ2):
(1) the effectiveness of baseline SD prompting in generating aligned distributions is model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al. (2024); (2) calibration increases opinion alignment in aggregate across models, datasets, and elicitation methods by an average of 16.3%; and (3) calibration reduces variance within and across settings.
4.1 Does SD prompting improve alignment?
To study RQ1, we compare SD vs. base prompts. What is the effect of adding SD information on opinion alignment? Note that while prior work studies this question only for majority-voted responses Beck et al. (2024); Hu and Collier (2024), we instead look at the effect of adding SD information when evaluating distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets. The results for all models are shown in Appendix Table 5.
For both base and SD prompts, LLM-generated distributions are generally best with verbalized elicitation (Table 1), though with exceptions (e.g., OLMo-Instruct on the OQA dataset). As shown in the table, prompting with SD information does not necessarily increase opinion alignment with human responses. In fact, it is often comparable or even lower than prompting without any SD information. The degree of effectiveness in creating distributions that align with human opinion appears to be model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al. (2024).
4.2 Can we calibrate LLM distributions?
We have seen that prompting LLMs with SD information does not consistently increase opinion alignment with human responses. Since LLM-generated distributions have fairly low opinion alignment, with RQ2, we question whether we can calibrate these distributions to better align them with human response distributions on these survey datasets.
The results after regression, compared to pre-regression results, are shown in Table 1, with results for all models in Appendix Table 5. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, and by an average of 16.3%. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment decreases by an average of 4.1%. These results suggest that LLM-generated distributions for predicting human responses are somewhat uncalibrated, and a simple supervised regression can lead to higher opinion alignment.
After calibration, opinion alignment in each setting also has much lower variance. Calibrated distributions tend to have a much lower standard deviation; standard deviation is lower in 87.2% of dataset-LLM-elicitation settings, and is on average 1.62 times lower. That lower standard deviation leads to higher opinion alignment provides more evidence that LLMs might be exaggerating differences in opinion distributions between demographic groups Cheng et al. (2023a, b), and that calibration could be helpful in mitigating this.
In addition, we find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. We see that the standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. While prior work Beck et al. (2024); Hu and Collier (2024) and our own previous results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly. This means that with calibration, the choice of any particular model becomes less important.
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| World region (WGM) | Aus/NZ | 82.3 | 79.7 | 94.7 | 83.5 | 88.2 | 86.6 | 73.5 | 79.1 | 89.1 | 84.2 | 85.6 | 82.6 |
| Central Africa | 85.5 | 87.7 | 71.2 | 81.2 | 78.1 | 78.6 | 38.0 | 79.0 | 75.6 | 79.2 | 69.7 | 81.1 | |
| Cent. America & Mex. | 89.2 | 90.0 | 74.8 | 85.1 | 81.7 | 82.8 | 46.4 | 84.7 | 78.5 | 81.8 | 74.1 | 84.9 | |
| Central Asia | 82.5 | 80.2 | 88.9 | 82.8 | 87.7 | 85.3 | 75.5 | 79.9 | 88.8 | 84.4 | 84.7 | 82.5 | |
| East Asia | 88.4 | 85.7 | 86.8 | 85.7 | 88.7 | 88.6 | 61.6 | 82.0 | 86.4 | 82.8 | 82.4 | 85.0 | |
| Eastern Africa | 88.2 | 88.4 | 80.0 | 87.7 | 88.0 | 87.3 | 63.3 | 87.7 | 86.0 | 86.2 | 81.1 | 87.5 | |
| Eastern Europe | 93.0 | 88.4 | 85.4 | 88.1 | 89.6 | 89.4 | 54.6 | 81.9 | 89.9 | 88.5 | 82.5 | 87.3 | |
| Middle East | 91.1 | 87.9 | 87.7 | 92.9 | 92.5 | 91.7 | 59.0 | 85.1 | 91.4 | 88.8 | 84.3 | 89.3 | |
| North Africa | 90.0 | 87.5 | 81.6 | 87.8 | 86.9 | 87.2 | 53.0 | 81.4 | 85.1 | 84.1 | 79.3 | 85.6 | |
| Northern America | 83.2 | 80.9 | 91.7 | 85.0 | 88.2 | 87.0 | 68.4 | 82.4 | 92.4 | 88.8 | 84.8 | 84.8 | |
| Northern Europe | 82.8 | 80.6 | 94.7 | 84.8 | 89.6 | 87.6 | 72.4 | 81.0 | 91.3 | 86.5 | 86.2 | 84.1 | |
| South America | 88.9 | 89.5 | 74.0 | 84.4 | 80.9 | 82.0 | 45.1 | 82.5 | 77.7 | 81.0 | 73.3 | 83.9 | |
| South Asia | 84.2 | 84.0 | 86.7 | 84.8 | 88.9 | 87.3 | 70.5 | 83.7 | 87.0 | 83.3 | 83.5 | 84.6 | |
| Southeast Asia | 82.1 | 82.5 | 79.1 | 80.2 | 86.3 | 84.6 | 70.6 | 81.9 | 84.2 | 81.7 | 80.5 | 82.2 | |
| Southern Africa | 86.4 | 90.1 | 74.2 | 84.4 | 82.4 | 82.8 | 49.1 | 88.8 | 79.7 | 83.3 | 74.4 | 85.9 | |
| Southern Europe | 93.2 | 89.1 | 84.5 | 88.7 | 90.0 | 89.7 | 53.4 | 81.7 | 88.5 | 88.2 | 81.9 | 87.5 | |
| Western Africa | 90.7 | 90.4 | 81.0 | 91.6 | 88.1 | 88.2 | 57.7 | 90.3 | 85.7 | 86.4 | 80.6 | 89.4 | |
| Western Europe | 83.2 | 80.6 | 93.0 | 86.4 | 90.3 | 88.5 | 71.6 | 84.2 | 93.5 | 88.7 | 86.3 | 85.7 | |
| All Demographics (WGM) | 89.3 | 87.7 | 84.8 | 89.0 | 89.4 | 88.9 | 59.8 | 85.1 | 88.2 | 87.1 | 82.3 | 87.5 | |
| P.I.(OQA) | Very conservative | 85.1 | 84.8 | 77.4 | 79.0 | 76.1 | 77.9 | 63.8 | 77.5 | 81.7 | 81.0 | 76.8 | 80.0 |
| Conservative | 89.0 | 89.0 | 82.4 | 83.7 | 81.3 | 82.9 | 64.4 | 79.2 | 85.7 | 85.1 | 80.6 | 84.0 | |
| Moderate | 92.3 | 92.3 | 87.9 | 88.9 | 85.7 | 87.0 | 65.0 | 79.5 | 88.5 | 88.0 | 83.9 | 87.1 | |
| Liberal | 91.6 | 92.1 | 89.5 | 89.3 | 85.5 | 85.9 | 63.1 | 77.8 | 88.6 | 88.9 | 83.7 | 86.8 | |
| Very liberal | 87.8 | 88.3 | 86.5 | 85.9 | 83.3 | 83.3 | 61.1 | 76.1 | 84.8 | 85.0 | 80.7 | 83.7 | |
| All Demographics (OQA) | 91.7 | 91.7 | 86.8 | 87.9 | 85.0 | 86.2 | 65.4 | 79.9 | 88.4 | 87.9 | 83.5 | 86.7 | |
| I.(WVS) | High | 77.1 | 82.2 | 66.0 | 82.2 | 74.3 | 82.2 | 87.6 | 89.9 | 75.9 | 79.8 | 76.2 | 83.3 |
| Middle | 75.2 | 80.2 | 64.7 | 80.2 | 72.7 | 80.2 | 86.8 | 86.6 | 73.6 | 77.8 | 74.6 | 81.0 | |
| Low | 77.4 | 82.4 | 66.8 | 82.4 | 75.0 | 82.4 | 87.8 | 88.5 | 76.1 | 80.0 | 76.6 | 83.1 | |
| All Demographics (WVS) | 75.2 | 80.3 | 64.6 | 80.3 | 72.8 | 80.3 | 86.6 | 86.5 | 73.9 | 77.9 | 74.6 | 81.1 | |
5 How does alignment vary across SDs?
So far, we have studied how well LLMs can predict responses from SD groups aggregated across all groups, with the goal of comparing different models and distribution elicitation methods. However, a particular model or method might be more aligned with some sociodemographics over others. With RQ3, we seek to understand the differences at the more granular demographic level.
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it performed the best with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. We choose 5 out of 15 LLMs; the most recent/powerful from each model family: Claude 3.5 v2, Llama 3.2 90B, Mistral large, OLMo-2 7B Instruct, and Qwen 2.5 72B. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 2. See Appendix A.4 for alignment for other demographics.
Findings (RQ3):
(1) Calibration produces distributions more aligned with some demographics over others; (2) Claude models are more highly aligned with OQA dataset SD groups; (3) different models are better aligned with specific demographics.
5.1 How does calibration affect alignment?
First, we compare opinion alignment between base and calibrated distributions. We use the regression models trained on all sociodemographic groups and study how such aggregate regression models might calibrate individual sociodemographics. We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more aligned for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. As expected, calibration increases alignment for some demographics and reduces alignment for others. We find that it tends to increase alignment of SD groups that were less aligned with base-prompted distributions, such as Central Africa, Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment is decreased for very high aligned demographic groups, such as Aus/NZ and Northern Europe with Llama 90B.
5.2 How does alignment vary across LLMs?
Next, we compare the opinion alignment between different LLMs. For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East. Interestingly, Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data. For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen. In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. This is discussed further in Section 7. All models have lowest alignment with middle income populations. This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income. In other words, income might be less of an important factor in determining the answer to the survey questions for those with middle income.
6 Does supervision work with less data?
Although we used 80% of our data as supervision for our regression models to calibrate the LLM-generated distributions, we study how opinion alignment varies with smaller amounts of supervised data to explore RQ4.
Findings (RQ4):
(1) As few as 5 training examples to the regression model can suffice to achieve close to minimal MSE; (2) on average, degradation in alignment for individual demographics is close to zero at 5 examples; (3) demographics with the largest opinion alignment degradation is dependent on base vs. SD prompts; and (4) regression models generalize fairly well to unseen datasets.
6.1 How much supervised data do you need?
We evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset. We average over 10 different random samples for each training data size. We find that as few as 5 examples (with 4 answer choices per example) can be enough to achieve close to the minimal MSE for any particular dataset-LLM-elicitation setting. We plot Mean Squared Error (MSE) over training data size in Appendix Figure 4 showing that MSE usually converges at 5 examples, though this is model and dataset dependent.
6.2 How are individual SD groups affected?
How does using a small set of random examples affect alignment of individual demographics? Although degradation on average is close to zero, individual demographics are affected differently. Table 3 shows those most affected and the degradation amount for both base and SD-prompted verbalized distributions. We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant. Looking at absolute values, using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
| Category | Demographic | ||||
| Base | Region | Southeast Asia | 81.0 | 79.4 | 1.6 |
| Education | Less than HS | 83.1 | 81.6 | 1.5 | |
| Race | Black | 86.6 | 85.1 | 1.5 | |
| Race | Hispanic | 87.2 | 85.9 | 1.4 | |
| Pol. Party | Democrat | 84.7 | 83.3 | 1.4 | |
| SD | Region | Northern Europe | 87.1 | 84.6 | 2.4 |
| Region | Western Europe | 87.1 | 84.9 | 2.3 | |
| Region | Aus/NZ | 84.3 | 82.1 | 2.2 | |
| Education | Tertiary | 88.6 | 86.5 | 2.1 | |
| Employment | Unemployed | 86.9 | 84.9 | 2.0 |
6.3 Do models generalize out-of-domain?
We also study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset. We train regression models on two of our three datasets and evaluate on the held-out dataset. Results are shown in Appendix A.5. As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset. Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM. This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
7 Does post-training impact alignment?
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + SFT, base + SFT + DPO, and the instruct model (base + SFT + DPO + RLVR). We study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it generally is best aligned.
Findings (RQ5):
Calibrated alignment is similar both before and after each phase of post-training, though pre-calibrated alignment decreases for 2/3 datasets when post-training is added.
7.1 What is the effect of calibration?
Figure 3 plots opinion alignment for all OLMo models. Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets. OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to affect opinion alignment much less.
7.2 What is the effect of post-training?
Changes in alignment after various post-training methods is dataset-dependent. For WGM, alignment of the base distribution falls significantly after post-training, though alignment for the SD distributions slightly increase with RLVR. Calibration alignment stays consistent after post-training. On OQA, alignment decreases with SFT, but stays about the same when adding other post-training. This is true of base, SD, and calibrated distributions, though calibration alignment decreases the least. On WVS however, alignment increases after post-training methods are added, and alignment changes less with calibration.
We suspect the increase in performance on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS. The drop in performance for uncalibrated distributions on WGM and OQA might indicate that OLMO’s post-training data is less aligned with some of the populations of these datasets.
8 Suggestions for practitioners
One must acknowledge risks and and exercise caution in using LLMs to approximate responses from diverse SD groups, since fallible LLMs may naturally exhibit random errors and systematic biases in their outputs. That said, the ease, speed, and cost of automated approximation may motivate some practitioners in some settings to use LLMs to approximate human responses, provided the predictive accuracy is “good enough” for their needs. Whether this is so will depend greatly on the varying accuracy needs of different real-world uses cases, as well as predictive accuracy relative to the SD groups of interest to each use case. We also note that such approximation may represent only an initial starting point, providing an initial guess of the data while awaiting for the real human response data to more slowly trickle in from participants.
In general, we suggest that practitioners: (1) evaluate distributions rather than a single response when evaluating alignment with human responses, and (2) calibrate LLM-generated distributions. As our results shown, calibration can be effective with as few as 5 supervised examples, though degradation varies different SD groups, so one must acknowledge the tradeoffs between risk, cost and quality in performing calibration with less supervision. We provide a summary of our results for the five models shown in Table 1.
8.1 Which methods and models work best?
As shown in Table 1, although the best distribution elicitation method is LLM and dataset dependent, verbalized elicitation leads to the best aligned distributions in the majority of settings and has a higher average across settings. Although this is true both pre-calibration and post-calibration, all methods are much more closely aligned post-calibration: without calibration, verbalized elicitation is the most aligned in 90% of settings, and with calibration it is the most aligned in 63% of settings. Adding sociodemographic information does not make LLM distributions more aligned with human distributions consistently; SD verbalized distributions are more aligned than their base verbalized counterparts in 67% of settings pre-calibration and 60% of settings post-calibration.
Overall, using verbalized elicitation (followed by calibration) might be best to obtain the most aligned distributions. However, we note that after calibration, the alignment of distributions produced by all the methods are fairly close (within 2% of each other). We also find that prompting with SD information is not as important, though might lead to slightly more aligned distributions in aggregate.
The most aligned models are much more dataset dependent. Claude 3.5 v2 is most aligned for WGM and OQA pre-calibration (with and without SD information). Post-calibration, Claude is most aligned for OQA, while Llama 3.2 90B is most aligned for WGM. For WVS, OLMo 7B Instruct is most aligned (both pre/post calibration).
9 Discussion
Essentializing sociodemographics. We note that by prompting LLMs with SD information – and evaluating their generated distributions against human participants with those sociodemographics– we are essentializing the sociodemographic to the identities of the survey respondents and their opinions Wang et al. (2024). However, practitioners still seek to understand opinions of specific populations (e.g., for election polling) where an aligned LLM might be beneficial for faster iteration on survey questions. Although sociodemographics are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023).
The effect of SD prompting. While our SD information provides valuable context about us, it alone does not determine our opinions, and it is fallacious to assume otherwise. As we have seen, prompting with SD information on its own also does not necessarily lead to more aligned distributions. Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al. (2023); Do et al. (2025). One could make personalized predictions for a sample of a group’s individuals members Gordon et al. (2022) rather than a single distributive prediction for the group. Multimodal prompts might also add SD information implicitly, though this might come with additional biases.
LLM distributions are uncalibrated. We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated – simple regression allows us to scale distributions to be much more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others. We also find that calibration reduces variance, both across examples and across settings. Since calibrated distributions generally have higher performance, this might provide evidence for LLMs exaggerating differences between sociodemographic groups; exaggerations that are reduced with calibration.
Calibration on individual SD groups. We do not find many significant differences between calibrated/uncalibrated distributions at the individual SD group level. Calibrating responses from each individual SD would likely improve alignment for that particular SD, which we do not do here. However, we caution efficacy may vary by demographic; some demographic traits might be more relevant to the response distributions of certain questions (e.g., urban vs. rural areas might affect gun rights questions more than marital status).
The dataset, LLM, and distribution elicitation method matter. As prior work has found when evaluating SD prompting on human majority responses Beck et al. (2024); Hu and Collier (2024), in this work we find that opinion alignment is dataset, LLM, and distribution elicitation method dependent. However, we find that verbalized elicitation tends to elicit probabilities that are most aligned with human distributions in the most settings, both before and after calibration. We find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is dataset-dependent.
10 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques. Our work also offers guidance for those using LLMs to predict human responses in practice, and our benchmark can enable and stimulate future research.
11 Limitations
Prompting methods. We do not study chain of thought or other prompting methods which might increase overall performance. We do not experiment with varied temperatures for our generations. For elicitation methods, though a larger might improve generated distributions, our choice of is reasonable given inference costs. We use for verbalized elicitation (as opposed to ) since we found little variance between runs. We also leave in-context learning or fine-tuning Orlikowski et al. (2025) with human responses to future work.
Logit-based elicitation methods. We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs. That said, we do also report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.3). We find that using log probabilities leads to the most aligned distributions in 10/24 or 42% of settings with calibration. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied. We only studied a subset of the demographics available in the survey datasets, and each demographic individually. Future work might study a larger subset of demographics, as well as intersectional demographics.
Regression. Our regression predicts each answer choice individually and then normalizes the predicted answers. We also optimize regression for MSE instead of the alignment metric. We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better. Future work might also ensemble multiple elicitation methods.
Ethics Statement
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for social good – for practitioners, NLP researchers, or social science researchers seeking to accurately approximate human responses for survey development, annotation tasks, or social science studies – this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
We also note that evaluating LLM responses against human responses of particular sociodemographic groups often assumes that everyone in a particular demographic group thinks the same way, and essentializes the demographic to an individual’s identity. In this work, we try to address this by evaluating against the distribution of human responses for a particular sociodemographic group. By doing so, we hope to better align LLMs with various sociodemographic groups in a distributionally pluralistic manner, rather than assume that all members of the group would answer the same way.
Acknowledgments
We thank Sooyong Lee for early contributions that did not make it into the ultimate paper. This research was supported in part by Cisco, Good Systems222https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), and a grant from Open Philanthropy. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix A Additional Results
A.1 Alignment results for all models
Results are in Table 5.
A.2 Minimal Supervision
Results are in Figure 4.
A.3 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024b) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 4.
| Model | Base prompt | SD prompt | |||
| WGM | Llama-3-70B | 73.1 | 86.2 | 67.5 | 86.9 |
| Llama-3.1-70B | 74.0 | 83.2 | 70.4 | 88.1 | |
| Llama-3.2-1B | 83.6 | 88.6 | 83.0 | 85.1 | |
| Llama-3.2-11B | 84.0 | 85.0 | 82.2 | 84.4 | |
| Average | 78.7 | 85.7 | 75.8 | 86.1 | |
| Std Dev | 5.9 | 2.3 | 8.0 | 1.7 | |
| OQA | Llama-3-70B | 75.4 | 85.2 | 72.3 | 88.1 |
| Llama-3.1-70B | 78.9 | 87.3 | 75.7 | 87.4 | |
| Llama-3.2-1B | 88.3 | 88.4 | 87.3 | 86.5 | |
| Llama-3.2-11B | 82.6 | 86.3 | 88.3 | 89.6 | |
| Average | 81.3 | 86.8 | 80.9 | 87.9 | |
| Std Dev | 5.5 | 1.4 | 8.1 | 1.3 | |
| WVS | Llama-3-70B | 54.1 | 82.8 | 51.1 | 83.5 |
| Llama-3.1-70B | 43.5 | 83.3 | 49.1 | 83.0 | |
| Llama-3.2-1B | 83.9 | 86.4 | 85.1 | 84.8 | |
| Llama-3.2-11B | 74.8 | 81.1 | 72.9 | 85.5 | |
| Average | 64.1 | 83.4 | 64.5 | 84.2 | |
| Std Dev | 18.6 | 2.2 | 17.4 | 1.1 | |
| Model | Base prompt | Sociodemographic prompt | |||||||||||
| WGM | OLMo-2-7B-Base | 65.6 | 80.1 | 79.2 | 81.8 | 82.2 | 83.9 | 74.8 | 81.9 | 71.0 | 81.8 | 74.9 | 82.9 |
| OLMo-2-7B-SFT | 80.6 | 81.8 | 70.3 | 81.8 | 68.6 | 86.3 | 75.6 | 82.6 | 70.2 | 82.1 | 65.4 | 85.0 | |
| OLMo-2-7B-DPO | 71.1 | 77.4 | 63.2 | 82.1 | 63.9 | 86.2 | 72.2 | 83.0 | 68.7 | 82.4 | 63.8 | 84.9 | |
| OLMo-2-7B-Instruct | 59.6 | 81.5 | 67.7 | 80.4 | 59.8 | 85.1 | 62.3 | 82.5 | 64.5 | 82.7 | 69.8 | 84.6 | |
| Llama-3-70B | 64.9 | 85.3 | 73.7 | 83.8 | 84.5 | 88.8 | 66.5 | 86.0 | 61.7 | 84.7 | 84.4 | 88.3 | |
| Llama-3.1-70B | 73.2 | 86.7 | 70.3 | 86.5 | 80.6 | 86.0 | 68.5 | 85.9 | 66.1 | 85.9 | 85.2 | 89.4 | |
| Llama-3.2-1B | 81.8 | 81.8 | 81.8 | 81.8 | – | – | 61.9 | 82.5 | 46.7 | 81.2 | 61.4 | 86.4 | |
| Llama-3.2-11B | 73.7 | 87.7 | 73.2 | 83.5 | 73.6 | 84.5 | 71.8 | 85.3 | 72.0 | 84.7 | 71.7 | 79.2 | |
| Llama-3.2-90B | 68.1 | 84.6 | 73.0 | 86.2 | 84.8 | 89.0 | 70.6 | 86.0 | 67.6 | 86.1 | 85.0 | 89.8 | |
| Qwen-2.5-72B | 57.7 | 83.1 | 53.3 | 83.2 | 88.2 | 87.1 | 66.0 | 84.1 | 63.4 | 85.2 | 89.1 | 89.4 | |
| Mistral-small | 61.6 | 83.9 | 61.6 | 84.6 | 88.9 | 88.8 | 67.5 | 86.6 | 64.4 | 85.8 | 89.0 | 89.1 | |
| Mistral-large | 62.4 | 84.7 | 72.0 | 83.9 | 89.4 | 88.9 | 68.4 | 84.7 | 63.0 | 84.9 | 87.3 | 88.4 | |
| Claude-3 | 57.8 | 81.1 | 50.0 | 81.8 | 84.8 | 86.7 | 64.9 | 80.6 | 52.9 | 81.7 | 88.4 | 88.1 | |
| Claude-3.5-v1 | 64.5 | 82.5 | 62.0 | 84.3 | 86.1 | 87.4 | 65.8 | 84.6 | 62.3 | 84.5 | 86.3 | 88.3 | |
| Claude-3.5-v2 | 66.2 | 85.2 | 59.5 | 85.1 | 89.3 | 87.7 | 65.4 | 84.4 | 61.4 | 84.2 | 89.0 | 89.3 | |
| Average | 67.3 | 83.2 | 67.4 | 83.4 | 80.3 | 86.9 | 68.1 | 84.0 | 63.7 | 83.9 | 79.4 | 86.9 | |
| Std Dev | 7.5 | 2.7 | 9.0 | 1.8 | 9.9 | 1.7 | 4.1 | 1.8 | 6.7 | 1.7 | 10.3 | 3.0 | |
| OQA | OLMo-2-7B-Base | 85.0 | 89.2 | 82.6 | 88.4 | 82.1 | 81.8 | 80.8 | 88.3 | 77.2 | 88.1 | 83.8 | 87.6 |
| OLMo-2-7B-SFT | 78.8 | 88.0 | 79.4 | 86.8 | 68.7 | 81.6 | 80.7 | 87.9 | 77.8 | 87.2 | 72.6 | 83.0 | |
| OLMo-2-7B-DPO | 79.3 | 88.6 | 81.5 | 86.7 | 69.3 | 82.8 | 82.1 | 88.4 | 78.7 | 87.7 | 72.3 | 82.7 | |
| OLMo-2-7B-Instruct | 72.6 | 89.0 | 72.3 | 88.4 | 65.4 | 79.9 | 72.8 | 88.4 | 70.5 | 88.6 | 68.5 | 81.5 | |
| Llama-3-70B | 72.2 | 87.5 | 76.8 | 89.1 | 81.8 | 85.0 | 74.4 | 88.9 | 70.1 | 88.9 | 79.5 | 83.7 | |
| Llama-3.1-70B | 73.5 | 88.8 | 70.5 | 87.7 | 83.8 | 86.4 | 76.4 | 89.5 | 71.7 | 89.4 | 83.6 | 86.6 | |
| Llama-3.2-1B | 82.3 | 88.4 | 88.4 | 88.4 | – | – | 72.6 | 89.0 | 70.2 | 88.6 | 83.6 | 84.9 | |
| Llama-3.2-11B | 84.3 | 89.4 | 75.5 | 86.4 | 65.8 | 77.2 | 75.7 | 87.2 | 74.8 | 87.0 | 83.0 | 86.9 | |
| Llama-3.2-90B | 79.3 | 89.6 | 75.3 | 89.4 | 86.8 | 87.9 | 79.2 | 90.1 | 76.1 | 90.0 | 83.4 | 85.9 | |
| Qwen-2.5-72B | 73.9 | 89.7 | 67.0 | 88.8 | 88.4 | 87.9 | 74.4 | 90.0 | 71.3 | 89.5 | 89.2 | 88.6 | |
| Mistral-small | 77.3 | 91.6 | 73.6 | 90.5 | 86.9 | 87.7 | 75.1 | 90.4 | 71.6 | 90.3 | 87.8 | 89.0 | |
| Mistral-large | 79.5 | 89.9 | 75.3 | 88.3 | 85.0 | 86.2 | 75.8 | 89.2 | 72.4 | 89.5 | 83.8 | 84.7 | |
| Claude-3 | 73.5 | 89.9 | 66.4 | 90.3 | 87.6 | 86.5 | 76.4 | 90.6 | 72.6 | 90.3 | 89.3 | 88.6 | |
| Claude-3.5-v1 | 70.5 | 86.9 | 69.6 | 88.6 | 89.4 | 89.0 | 76.1 | 90.3 | 72.8 | 89.8 | 91.0 | 90.8 | |
| Claude-3.5-v2 | 70.5 | 88.8 | 72.5 | 90.4 | 91.7 | 91.7 | 76.1 | 89.9 | 73.3 | 89.6 | 91.9 | 91.6 | |
| Average | 76.8 | 89.0 | 75.1 | 88.6 | 80.9 | 85.1 | 76.6 | 89.2 | 73.4 | 89.0 | 82.9 | 86.4 | |
| Std Dev | 4.8 | 1.1 | 6.0 | 1.3 | 9.4 | 4.0 | 2.9 | 1.0 | 2.8 | 1.1 | 7.0 | 3.0 | |
| WVS | OLMo-2-7B-Base | 74.5 | 80.3 | 70.7 | 80.1 | 84.5 | 80.3 | 78.2 | 81.4 | 78.2 | 81.5 | 80.8 | 81.9 |
| OLMo-2-7B-SFT | 69.7 | 79.2 | 68.1 | 79.1 | 87.2 | 87.2 | 72.9 | 80.6 | 73.1 | 80.7 | 84.9 | 89.0 | |
| OLMo-2-7B-DPO | 66.2 | 80.3 | 70.5 | 79.8 | 89.1 | 83.9 | 70.6 | 80.9 | 70.2 | 80.7 | 82.5 | 85.6 | |
| OLMo-2-7B-Instruct | 75.3 | 80.3 | 74.9 | 80.3 | 86.6 | 86.5 | 58.3 | 79.2 | 60.0 | 77.2 | 86.0 | 89.8 | |
| Llama-3-70B | 59.7 | 80.3 | 61.7 | 80.6 | 81.3 | 83.0 | 67.9 | 80.8 | 62.8 | 80.5 | 78.1 | 82.1 | |
| Llama-3.1-70B | 61.9 | 81.5 | 56.4 | 82.1 | 65.2 | 80.3 | 63.6 | 80.8 | 57.6 | 81.3 | 68.3 | 82.8 | |
| Llama-3.2-1B | 80.3 | 80.3 | 80.3 | 80.3 | – | – | 68.7 | 82.8 | 66.6 | 78.9 | – | – | |
| Llama-3.2-11B | 75.4 | 82.0 | 76.0 | 84.8 | 82.7 | 84.0 | 78.2 | 82.2 | 78.0 | 81.2 | 66.5 | 87.0 | |
| Llama-3.2-90B | 61.6 | 79.9 | 59.1 | 80.3 | 64.6 | 80.3 | 62.1 | 80.8 | 59.5 | 81.5 | 67.7 | 82.7 | |
| Qwen-2.5-72B | 39.6 | 82.0 | 42.4 | 82.3 | 74.0 | 77.9 | 49.1 | 82.4 | 49.4 | 81.5 | 73.4 | 81.7 | |
| Mistral-small | 42.9 | 80.3 | 46.3 | 80.3 | 75.0 | 80.3 | 48.0 | 80.4 | 46.4 | 80.4 | 68.6 | 81.7 | |
| Mistral-large | 48.8 | 80.3 | 44.0 | 82.2 | 72.8 | 80.3 | 54.3 | 80.4 | 51.5 | 80.3 | 76.6 | 83.8 | |
| Claude-3 | 47.3 | 80.3 | 55.7 | 80.3 | 74.1 | 80.3 | 55.3 | 80.2 | 57.3 | 79.8 | 73.2 | 81.7 | |
| Claude-3.5-v1 | 44.3 | 80.3 | 49.6 | 80.3 | 75.7 | 80.3 | 58.7 | 80.3 | 54.8 | 80.3 | 73.3 | 81.6 | |
| Claude-3.5-v2 | 46.5 | 80.3 | 51.0 | 80.3 | 75.2 | 80.3 | 61.0 | 80.4 | 56.8 | 80.4 | 75.6 | 81.7 | |
| Average | 59.6 | 80.5 | 60.4 | 80.9 | 77.7 | 81.8 | 63.1 | 80.9 | 61.5 | 80.4 | 75.4 | 83.8 | |
| Std Dev | 13.8 | 0.7 | 12.4 | 1.4 | 7.8 | 2.7 | 9.5 | 0.9 | 9.9 | 1.1 | 6.4 | 2.9 | |
A.4 Individual Sociodemographic Results
Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 6, 7, and 8 respectively.
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| Age | 15-29 | 95.2 | 92.7 | 85.0 | 94.1 | 92.9 | 92.5 | 56.6 | 88.6 | 90.2 | 90.7 | 84.0 | 91.7 |
| 30-49 | 92.7 | 90.7 | 85.7 | 93.1 | 93.1 | 92.7 | 58.9 | 87.0 | 90.6 | 89.8 | 84.2 | 90.7 | |
| 50+ | 88.6 | 87.5 | 87.9 | 91.0 | 91.4 | 90.7 | 63.7 | 85.9 | 90.6 | 87.1 | 84.4 | 88.4 | |
| Edu. | primary | 83.9 | 83.8 | 78.0 | 86.3 | 84.6 | 84.0 | 62.1 | 88.6 | 82.5 | 82.6 | 78.2 | 85.1 |
| secondary | 95.1 | 91.2 | 86.5 | 91.8 | 92.1 | 92.2 | 57.0 | 85.8 | 91.1 | 90.5 | 84.4 | 90.3 | |
| tertiary | 86.2 | 82.7 | 88.6 | 88.4 | 90.4 | 89.5 | 63.8 | 86.2 | 94.6 | 90.6 | 84.7 | 87.5 | |
| Employment | full time for employer | 92.3 | 88.5 | 89.1 | 91.1 | 93.4 | 92.7 | 60.6 | 86.0 | 93.9 | 90.8 | 85.9 | 89.8 |
| full time for self | 91.0 | 89.7 | 84.3 | 91.3 | 91.5 | 91.2 | 60.0 | 88.2 | 89.3 | 88.8 | 83.2 | 89.8 | |
| part time (no full time) | 91.2 | 89.9 | 86.0 | 93.2 | 92.4 | 92.1 | 60.9 | 88.1 | 90.2 | 89.2 | 84.1 | 90.5 | |
| part time (full time) | 92.1 | 90.6 | 83.1 | 93.0 | 90.3 | 90.4 | 58.5 | 89.7 | 88.0 | 87.9 | 82.4 | 90.3 | |
| out of work force | 91.1 | 89.6 | 85.6 | 92.1 | 91.5 | 91.1 | 60.5 | 86.7 | 89.7 | 88.6 | 83.7 | 89.6 | |
| unemployed | 92.9 | 92.7 | 80.3 | 91.4 | 88.1 | 88.8 | 51.6 | 87.4 | 85.0 | 87.5 | 79.6 | 89.6 | |
| Sex | female | 92.2 | 90.5 | 84.6 | 92.2 | 91.8 | 91.6 | 59.0 | 87.4 | 89.3 | 88.8 | 83.4 | 90.1 |
| male | 91.9 | 89.5 | 87.8 | 93.4 | 93.8 | 92.6 | 60.6 | 86.8 | 92.5 | 90.5 | 85.3 | 90.6 | |
| Income | fourth 20% | 92.9 | 90.7 | 87.5 | 93.3 | 94.3 | 93.2 | 60.1 | 86.9 | 92.0 | 90.3 | 85.4 | 90.9 |
| middle 20% | 91.9 | 90.3 | 86.2 | 92.5 | 92.7 | 92.4 | 60.2 | 87.1 | 90.7 | 89.4 | 84.3 | 90.3 | |
| poorest 20% | 90.2 | 88.8 | 82.5 | 91.3 | 89.2 | 89.1 | 59.0 | 87.9 | 86.9 | 87.2 | 81.6 | 88.9 | |
| second 20% | 91.0 | 89.5 | 85.0 | 91.7 | 91.4 | 91.1 | 60.4 | 87.3 | 89.4 | 88.4 | 83.4 | 89.6 | |
| top 20% | 92.5 | 89.1 | 87.5 | 92.5 | 93.6 | 92.7 | 59.5 | 86.8 | 93.7 | 92.0 | 85.4 | 90.6 | |
| Area | city/suburb | 94.2 | 91.4 | 85.9 | 92.7 | 93.0 | 93.2 | 56.4 | 86.3 | 90.4 | 90.3 | 84.0 | 90.8 |
| rural/small town | 90.3 | 89.2 | 86.2 | 92.0 | 92.0 | 91.5 | 62.1 | 87.8 | 90.2 | 88.5 | 84.2 | 89.8 | |
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| Age | 18-29 | 92.5 | 92.6 | 87.5 | 88.8 | 86.9 | 88.0 | 68.9 | 83.2 | 89.7 | 89.2 | 85.1 | 88.4 |
| 30-49 | 92.6 | 92.5 | 86.5 | 87.8 | 85.9 | 87.1 | 66.6 | 81.5 | 89.2 | 88.7 | 84.2 | 87.5 | |
| 50-64 | 92.5 | 92.4 | 86.5 | 87.8 | 84.1 | 85.8 | 63.5 | 78.6 | 88.1 | 87.5 | 82.9 | 86.4 | |
| 65+ | 92.1 | 92.1 | 87.3 | 88.3 | 83.9 | 85.4 | 62.1 | 76.9 | 87.9 | 87.3 | 82.7 | 86.0 | |
| Edu. | Associate’s degree | 92.5 | 92.4 | 85.9 | 87.4 | 84.6 | 86.5 | 65.3 | 80.5 | 88.3 | 87.6 | 83.3 | 86.9 |
| College graduate | 92.1 | 92.0 | 87.6 | 88.6 | 85.7 | 87.0 | 65.0 | 79.8 | 88.6 | 88.0 | 83.8 | 87.1 | |
| High school graduate | 91.5 | 91.4 | 85.2 | 86.5 | 83.2 | 84.8 | 63.6 | 78.5 | 87.8 | 87.2 | 82.3 | 85.7 | |
| Less than high school | 88.8 | 88.3 | 84.4 | 85.5 | 82.7 | 84.0 | 69.2 | 81.3 | 87.0 | 86.5 | 82.4 | 85.1 | |
| Postgraduate | 92.2 | 92.1 | 87.3 | 88.4 | 85.4 | 86.3 | 64.0 | 78.7 | 88.2 | 87.6 | 83.4 | 86.6 | |
| Some college, no degree | 93.7 | 93.7 | 87.0 | 88.4 | 84.9 | 86.5 | 65.2 | 80.2 | 89.4 | 88.8 | 84.0 | 87.5 | |
| Region | Midwest | 92.3 | 92.3 | 86.2 | 87.6 | 84.7 | 86.2 | 64.2 | 79.3 | 88.3 | 87.7 | 83.1 | 86.6 |
| Northeast | 92.9 | 92.8 | 88.2 | 89.1 | 85.6 | 87.1 | 64.0 | 78.7 | 89.1 | 88.6 | 84.0 | 87.3 | |
| South | 92.9 | 92.8 | 87.1 | 88.3 | 84.9 | 86.6 | 65.1 | 79.9 | 88.9 | 88.4 | 83.8 | 87.2 | |
| West | 92.8 | 92.7 | 87.9 | 89.1 | 86.0 | 87.3 | 65.0 | 80.1 | 89.1 | 88.5 | 84.2 | 87.5 | |
| Income | $100,000 or more | 90.7 | 90.6 | 85.7 | 86.9 | 84.1 | 85.3 | 64.0 | 78.6 | 86.9 | 86.2 | 82.3 | 85.5 |
| $30,000-$50,000 | 93.2 | 93.1 | 86.3 | 87.6 | 84.9 | 86.5 | 65.1 | 80.0 | 89.4 | 88.8 | 83.8 | 87.2 | |
| $50,000-$75,000 | 93.2 | 93.2 | 87.1 | 88.3 | 85.4 | 86.8 | 64.9 | 79.7 | 89.1 | 88.5 | 83.9 | 87.3 | |
| $75,000-$100,000 | 93.0 | 92.8 | 87.9 | 89.2 | 86.1 | 87.5 | 65.6 | 80.4 | 89.4 | 88.8 | 84.4 | 87.7 | |
| Less than $30,000 | 93.9 | 93.7 | 88.4 | 89.8 | 86.1 | 87.7 | 65.7 | 80.7 | 90.2 | 89.5 | 84.9 | 88.3 | |
| Marital | Divorced | 94.3 | 94.1 | 86.8 | 88.3 | 85.0 | 86.8 | 63.7 | 78.8 | 89.5 | 88.7 | 83.9 | 87.3 |
| Married | 92.0 | 91.9 | 86.5 | 87.8 | 84.7 | 86.1 | 64.6 | 79.5 | 88.3 | 87.7 | 83.2 | 86.6 | |
| Never been married | 93.3 | 93.2 | 88.5 | 89.9 | 87.5 | 88.5 | 66.5 | 81.4 | 89.9 | 89.3 | 85.1 | 88.5 | |
| Separated | 93.8 | 93.6 | 89.2 | 90.3 | 86.4 | 87.6 | 65.4 | 79.4 | 90.1 | 89.7 | 85.0 | 88.1 | |
| Widowed | 90.1 | 90.0 | 87.5 | 88.4 | 82.9 | 84.1 | 61.5 | 75.8 | 86.1 | 85.5 | 81.6 | 84.8 | |
| Pol. Party | Democrat | 90.5 | 91.0 | 89.5 | 89.2 | 85.3 | 85.8 | 62.5 | 77.3 | 87.9 | 88.4 | 83.1 | 86.3 |
| Independent | 92.1 | 92.0 | 86.3 | 87.4 | 84.3 | 85.7 | 64.6 | 79.6 | 88.2 | 87.6 | 83.1 | 86.5 | |
| Other | 91.2 | 91.0 | 84.5 | 86.1 | 84.0 | 85.6 | 65.8 | 80.9 | 87.5 | 86.7 | 82.6 | 86.1 | |
| Republican | 88.3 | 88.2 | 81.9 | 83.2 | 80.6 | 82.2 | 64.7 | 79.5 | 85.4 | 84.9 | 80.2 | 83.6 | |
| Race | Asian | 93.9 | 93.8 | 88.3 | 89.8 | 89.8 | 90.8 | 73.1 | 86.9 | 92.6 | 92.0 | 87.5 | 90.7 |
| Black | 93.4 | 93.6 | 91.4 | 91.8 | 87.4 | 88.1 | 65.7 | 80.3 | 90.5 | 90.0 | 85.7 | 88.8 | |
| Hispanic | 94.3 | 94.3 | 90.3 | 91.2 | 87.6 | 88.6 | 67.1 | 81.9 | 91.6 | 91.1 | 86.2 | 89.4 | |
| Other | 93.1 | 92.5 | 86.5 | 88.3 | 85.9 | 87.7 | 67.2 | 81.6 | 90.4 | 89.5 | 84.6 | 87.9 | |
| White | 92.2 | 92.1 | 86.4 | 87.7 | 84.5 | 85.9 | 63.8 | 78.8 | 88.2 | 87.6 | 83.0 | 86.4 | |
| Religion | Agnostic | 89.5 | 89.5 | 85.1 | 86.4 | 82.9 | 84.1 | 63.3 | 77.4 | 85.7 | 84.9 | 81.3 | 84.5 |
| Atheist | 87.6 | 88.1 | 85.3 | 85.9 | 82.4 | 83.2 | 62.4 | 76.8 | 84.6 | 84.1 | 80.5 | 83.6 | |
| Buddhist | 90.1 | 90.2 | 90.2 | 89.9 | 86.4 | 86.5 | 66.4 | 81.0 | 88.5 | 88.5 | 84.3 | 87.2 | |
| Hindu | 88.3 | 88.6 | 86.3 | 85.9 | 86.9 | 86.4 | 73.0 | 85.2 | 87.2 | 87.5 | 84.3 | 86.7 | |
| Jewish | 93.3 | 93.6 | 87.7 | 88.7 | 87.2 | 88.2 | 66.8 | 80.7 | 89.9 | 89.5 | 85.0 | 88.1 | |
| Mormon | 87.5 | 87.2 | 82.6 | 84.0 | 82.9 | 84.2 | 67.8 | 81.4 | 85.8 | 85.2 | 81.3 | 84.4 | |
| Muslim | 91.9 | 91.6 | 91.2 | 92.3 | 89.4 | 90.7 | 69.9 | 83.9 | 91.4 | 90.7 | 86.8 | 89.8 | |
| Nothing in particular | 93.2 | 93.1 | 89.2 | 90.3 | 86.6 | 88.0 | 65.1 | 80.2 | 89.4 | 89.0 | 84.7 | 88.1 | |
| Orthodox | 93.4 | 93.1 | 88.5 | 89.9 | 88.2 | 89.3 | 69.9 | 84.1 | 91.8 | 91.2 | 86.4 | 89.5 | |
| Other | 92.7 | 92.5 | 86.1 | 87.6 | 84.0 | 85.6 | 63.4 | 78.5 | 89.5 | 88.9 | 83.1 | 86.6 | |
| Protestant | 91.4 | 91.4 | 85.1 | 86.3 | 83.3 | 85.0 | 63.9 | 78.8 | 87.3 | 86.8 | 82.2 | 85.7 | |
| Roman Catholic | 93.1 | 93.0 | 87.5 | 88.7 | 85.7 | 87.2 | 66.3 | 80.8 | 89.3 | 88.7 | 84.4 | 87.7 | |
| Sex | Female | 93.4 | 93.3 | 87.8 | 89.1 | 85.1 | 86.5 | 63.5 | 78.4 | 89.2 | 88.7 | 83.8 | 87.2 |
| Male | 91.5 | 91.4 | 86.1 | 87.6 | 85.4 | 86.6 | 66.2 | 81.1 | 88.1 | 87.4 | 83.5 | 86.8 | |
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| Age | 16-24 years | 76.4 | 81.6 | 65.5 | 81.6 | 73.9 | 81.6 | 86.2 | 88.3 | 75.3 | 79.2 | 75.5 | 82.5 |
| 25-34 years | 76.8 | 81.8 | 65.9 | 81.8 | 74.3 | 81.8 | 86.6 | 88.8 | 75.6 | 79.4 | 75.8 | 82.7 | |
| 35-44 years | 76.7 | 81.7 | 65.9 | 81.7 | 74.1 | 81.7 | 87.1 | 88.2 | 75.2 | 79.2 | 75.8 | 82.5 | |
| 45-54 years | 76.0 | 81.0 | 65.5 | 81.0 | 73.6 | 81.0 | 87.5 | 87.6 | 74.6 | 78.5 | 75.4 | 81.8 | |
| 55-64 years | 74.6 | 79.5 | 64.3 | 79.5 | 72.3 | 79.5 | 87.7 | 85.8 | 73.2 | 77.1 | 74.4 | 80.3 | |
| 65+ years | 72.9 | 77.9 | 62.8 | 77.9 | 70.8 | 77.9 | 87.5 | 84.6 | 71.6 | 75.5 | 73.1 | 78.8 | |
| Education | bachelor | 74.1 | 79.1 | 63.7 | 79.1 | 71.8 | 79.1 | 86.9 | 85.3 | 72.8 | 76.7 | 73.9 | 79.9 |
| doctoral | 73.3 | 78.8 | 62.6 | 78.8 | 71.0 | 78.8 | 86.0 | 85.7 | 72.5 | 76.4 | 73.1 | 79.7 | |
| early childhood | 78.7 | 83.7 | 68.1 | 83.7 | 76.4 | 83.7 | 90.2 | 91.3 | 77.4 | 81.2 | 78.2 | 84.7 | |
| lower secondary | 75.6 | 80.6 | 65.3 | 80.6 | 73.4 | 80.6 | 87.6 | 87.1 | 74.3 | 78.2 | 75.2 | 81.4 | |
| master | 74.3 | 79.5 | 63.6 | 79.5 | 72.0 | 79.5 | 86.8 | 87.0 | 73.2 | 77.1 | 74.0 | 80.5 | |
| post-secondary | 75.4 | 80.4 | 64.6 | 80.4 | 72.8 | 80.4 | 86.2 | 86.2 | 74.0 | 78.0 | 74.6 | 81.1 | |
| primary | 76.1 | 81.1 | 65.7 | 81.1 | 73.9 | 81.1 | 88.8 | 87.7 | 74.8 | 78.7 | 75.9 | 81.9 | |
| short-cycle tertiary | 73.7 | 78.7 | 63.2 | 78.7 | 71.4 | 78.7 | 85.9 | 85.2 | 72.3 | 76.2 | 73.3 | 79.5 | |
| upper secondary | 75.4 | 80.3 | 64.6 | 80.3 | 72.9 | 80.3 | 86.2 | 86.7 | 74.0 | 77.9 | 74.6 | 81.1 | |
| Employment | full time | 75.1 | 80.1 | 64.7 | 80.1 | 72.8 | 80.1 | 87.1 | 86.5 | 73.7 | 77.7 | 74.7 | 80.9 |
| housewife | 75.7 | 80.7 | 65.4 | 80.7 | 73.5 | 80.7 | 87.5 | 87.9 | 74.4 | 78.3 | 75.3 | 81.7 | |
| other | 74.8 | 79.8 | 64.7 | 79.8 | 72.7 | 79.8 | 87.1 | 83.9 | 73.6 | 77.4 | 74.6 | 80.1 | |
| part time | 76.8 | 81.8 | 66.2 | 81.8 | 74.3 | 81.8 | 87.1 | 88.2 | 75.3 | 79.4 | 75.9 | 82.6 | |
| retired/pensioned | 72.9 | 77.9 | 62.6 | 77.9 | 70.7 | 77.9 | 87.1 | 84.8 | 71.6 | 75.4 | 73.0 | 78.8 | |
| self-employed | 77.0 | 82.0 | 66.0 | 82.0 | 74.4 | 82.0 | 87.4 | 89.6 | 75.7 | 79.6 | 76.1 | 83.0 | |
| student | 75.6 | 80.8 | 64.9 | 80.8 | 73.1 | 80.8 | 86.1 | 87.0 | 74.4 | 78.4 | 74.8 | 81.6 | |
| unemployed | 78.0 | 83.0 | 67.0 | 83.0 | 75.4 | 83.0 | 86.9 | 89.0 | 76.8 | 80.6 | 76.8 | 83.7 | |
| Household size | 1 | 74.1 | 79.1 | 63.8 | 79.1 | 72.0 | 79.1 | 87.2 | 85.0 | 72.9 | 76.7 | 74.0 | 79.8 |
| 2 | 73.5 | 78.5 | 63.3 | 78.5 | 71.4 | 78.5 | 87.0 | 84.6 | 72.2 | 76.1 | 73.5 | 79.2 | |
| 3 | 75.0 | 80.0 | 64.5 | 80.0 | 72.6 | 80.0 | 87.1 | 85.9 | 73.4 | 77.5 | 74.5 | 80.7 | |
| 4 | 75.6 | 80.5 | 65.1 | 80.5 | 73.1 | 80.5 | 86.9 | 86.7 | 74.0 | 78.1 | 74.9 | 81.3 | |
| 5 | 76.5 | 81.4 | 65.5 | 81.4 | 73.9 | 81.4 | 87.0 | 88.8 | 75.2 | 79.0 | 75.6 | 82.4 | |
| 6 | 76.8 | 81.8 | 65.7 | 81.8 | 74.1 | 81.8 | 87.2 | 89.5 | 75.6 | 79.4 | 75.9 | 82.9 | |
| 7 persons or more | 78.9 | 86.5 | 66.1 | 86.5 | 74.6 | 86.5 | 80.6 | 86.0 | 78.4 | 84.8 | 75.7 | 86.1 | |
| Imm. | Immigrant | 71.7 | 76.7 | 61.3 | 76.7 | 69.5 | 76.7 | 86.0 | 83.8 | 70.4 | 74.2 | 71.8 | 77.6 |
| Native | 76.1 | 81.1 | 65.5 | 81.1 | 73.5 | 81.1 | 87.0 | 87.6 | 74.6 | 78.7 | 75.3 | 81.9 | |
| Marital | Divorced | 74.2 | 79.2 | 63.9 | 79.2 | 72.0 | 79.2 | 86.7 | 84.8 | 72.9 | 76.8 | 73.9 | 79.8 |
| Living together | 74.1 | 79.2 | 63.5 | 79.2 | 71.7 | 79.2 | 83.8 | 82.7 | 72.6 | 76.7 | 73.1 | 79.4 | |
| Married | 76.2 | 81.2 | 65.6 | 81.2 | 73.6 | 81.2 | 87.4 | 88.3 | 74.7 | 78.8 | 75.5 | 82.1 | |
| Separated | 74.5 | 79.4 | 63.9 | 79.4 | 72.1 | 79.4 | 84.9 | 83.6 | 73.1 | 77.0 | 73.7 | 79.8 | |
| Single | 75.9 | 81.0 | 65.3 | 81.0 | 73.5 | 81.0 | 86.6 | 87.1 | 74.5 | 78.6 | 75.2 | 81.7 | |
| Widowed | 75.0 | 80.0 | 64.2 | 80.0 | 72.5 | 80.0 | 87.0 | 86.5 | 73.7 | 77.5 | 74.5 | 80.8 | |
| Religion | Buddhist | 74.8 | 79.8 | 63.8 | 79.8 | 72.1 | 79.8 | 85.8 | 86.1 | 73.5 | 77.3 | 74.0 | 80.6 |
| Hindu | 79.7 | 86.0 | 68.4 | 86.0 | 76.9 | 86.0 | 85.3 | 93.4 | 79.6 | 83.6 | 78.0 | 87.0 | |
| Jew | 76.3 | 81.5 | 65.2 | 81.5 | 73.6 | 81.5 | 88.0 | 87.9 | 75.2 | 79.0 | 75.7 | 82.3 | |
| Muslim | 76.9 | 81.8 | 66.1 | 81.8 | 74.4 | 81.8 | 89.5 | 90.5 | 75.6 | 79.4 | 76.5 | 83.1 | |
| Orthodox | 73.1 | 78.6 | 62.3 | 78.6 | 70.7 | 78.6 | 85.5 | 85.8 | 72.4 | 76.2 | 72.8 | 79.6 | |
| Other | 70.8 | 75.7 | 61.4 | 75.7 | 68.3 | 75.7 | 85.7 | 80.1 | 69.2 | 73.3 | 71.1 | 76.1 | |
| Other Christian | 71.6 | 76.6 | 61.5 | 76.6 | 69.5 | 76.6 | 84.0 | 80.6 | 70.3 | 74.2 | 71.4 | 76.9 | |
| Protestant | 74.3 | 79.3 | 63.7 | 79.3 | 71.9 | 79.3 | 85.3 | 83.9 | 72.9 | 76.9 | 73.6 | 79.7 | |
| Roman Catholic | 73.9 | 78.9 | 63.4 | 78.9 | 71.6 | 78.9 | 85.2 | 83.8 | 72.6 | 76.5 | 73.3 | 79.4 | |
| none | 70.6 | 75.6 | 61.0 | 75.6 | 68.5 | 75.6 | 84.9 | 81.1 | 69.3 | 73.2 | 70.9 | 76.2 | |
| Sex | Female | 75.2 | 80.2 | 64.3 | 80.2 | 72.5 | 80.2 | 86.1 | 86.5 | 73.8 | 77.7 | 74.4 | 81.0 |
| Male | 76.3 | 81.2 | 65.9 | 81.2 | 73.9 | 81.2 | 87.8 | 87.7 | 74.8 | 78.8 | 75.7 | 82.0 | |
A.5 Generalization results
| Model | Base prompt | Sociodemographic prompt | |||||||||||
| WGM | OLMo-2-7B-Base | 65.6 | 82.4 | 79.2 | 81.8 | 82.2 | 82.6 | 74.8 | 81.9 | 71.0 | 82.1 | 74.9 | 82.6 |
| OLMo-2-7B-SFT | 80.6 | 81.8 | 70.3 | 81.3 | 68.6 | 83.9 | 75.6 | 82.4 | 70.2 | 82.2 | 65.4 | 82.7 | |
| OLMo-2-7B-DPO | 71.1 | 81.7 | 63.2 | 82.1 | 63.9 | 83.1 | 72.2 | 82.7 | 68.7 | 82.4 | 63.8 | 83.0 | |
| OLMo-2-7B-Instruct | 59.6 | 82.0 | 67.7 | 81.8 | 59.8 | 82.8 | 62.3 | 81.7 | 64.5 | 82.2 | 69.8 | 84.4 | |
| Llama-3-70B | 64.9 | 83.6 | 73.7 | 81.8 | 84.5 | 86.5 | 66.5 | 84.6 | 61.7 | 84.1 | 84.4 | 87.1 | |
| Llama-3.1-70B | 73.2 | 82.7 | 70.3 | 84.4 | 80.6 | 81.8 | 68.5 | 85.2 | 66.1 | 81.8 | 85.2 | 87.9 | |
| Llama-3.2-1B | 81.8 | 81.8 | 81.8 | 81.8 | – | – | 61.9 | 82.3 | 46.7 | 80.7 | 61.4 | 86.4 | |
| Llama-3.2-11B | 73.7 | 85.9 | 73.2 | 81.4 | 73.6 | 82.1 | 71.8 | 84.5 | 72.0 | 84.1 | 71.7 | 81.1 | |
| Llama-3.2-90B | 68.1 | 84.1 | 73.0 | 84.4 | 84.8 | 86.2 | 70.6 | 84.7 | 67.6 | 84.4 | 85.0 | 86.4 | |
| Qwen-2.5-72B | 57.7 | 83.5 | 53.3 | 83.0 | 88.2 | 86.7 | 66.0 | 81.8 | 63.4 | 83.7 | 89.1 | 85.9 | |
| Mistral-small | 61.6 | 84.1 | 61.6 | 83.7 | 88.9 | 85.7 | 67.5 | 84.6 | 64.4 | 83.9 | 89.0 | 85.0 | |
| Mistral-large | 62.4 | 84.4 | 72.0 | 83.5 | 89.4 | 86.8 | 68.4 | 84.6 | 63.0 | 84.2 | 87.3 | 86.2 | |
| Claude-3 | 57.8 | 82.2 | 50.0 | 81.8 | 84.8 | 84.3 | 64.9 | 82.2 | 52.9 | 81.6 | 88.4 | 85.6 | |
| Claude-3.5-v1 | 64.5 | 83.9 | 62.0 | 83.5 | 86.1 | 84.9 | 65.8 | 84.7 | 62.3 | 84.5 | 86.3 | 88.1 | |
| Claude-3.5-v2 | 66.2 | 84.3 | 59.5 | 83.1 | 89.3 | 86.5 | 65.4 | 84.3 | 61.4 | 84.1 | 89.0 | 87.7 | |
| Average | 67.3 | 83.2 | 67.4 | 82.6 | 80.3 | 84.6 | 68.1 | 83.5 | 63.7 | 83.1 | 79.4 | 85.3 | |
| Std Dev | 7.5 | 1.2 | 9.0 | 1.1 | 9.9 | 1.8 | 4.1 | 1.3 | 6.7 | 1.2 | 10.3 | 2.2 | |
| OQA | OLMo-2-7B-Base | 85.0 | 89.1 | 82.6 | 88.4 | 82.1 | 88.0 | 80.8 | 88.3 | 77.2 | 88.3 | 83.8 | 88.4 |
| OLMo-2-7B-SFT | 78.8 | 88.4 | 79.4 | 88.4 | 68.7 | 83.4 | 80.7 | 88.3 | 77.8 | 88.4 | 72.6 | 84.2 | |
| OLMo-2-7B-DPO | 79.3 | 88.4 | 81.5 | 88.3 | 69.3 | 83.6 | 82.1 | 88.2 | 78.7 | 88.3 | 72.3 | 85.4 | |
| OLMo-2-7B-Instruct | 72.6 | 88.6 | 72.3 | 88.4 | 65.4 | 82.5 | 72.8 | 88.5 | 70.5 | 88.6 | 68.5 | 81.4 | |
| Llama-3-70B | 72.2 | 88.4 | 76.8 | 88.4 | 81.8 | 88.4 | 74.4 | 88.7 | 70.1 | 89.0 | 79.5 | 88.8 | |
| Llama-3.1-70B | 73.5 | 88.9 | 70.5 | 88.9 | 83.8 | 88.4 | 76.4 | 89.1 | 71.7 | 90.4 | 83.6 | 89.6 | |
| Llama-3.2-1B | 82.3 | 88.4 | 88.4 | 88.4 | – | – | 72.6 | 89.1 | 70.2 | 88.5 | 83.6 | 87.6 | |
| Llama-3.2-11B | 84.3 | 88.4 | 75.5 | 87.7 | 65.8 | 88.2 | 75.7 | 87.8 | 74.8 | 87.7 | 83.0 | 88.7 | |
| Llama-3.2-90B | 79.3 | 89.4 | 75.3 | 88.4 | 86.8 | 88.4 | 79.2 | 88.9 | 76.1 | 89.0 | 83.4 | 88.8 | |
| Qwen-2.5-72B | 73.9 | 88.4 | 67.0 | 88.5 | 88.4 | 88.9 | 74.4 | 88.4 | 71.3 | 88.4 | 89.2 | 89.1 | |
| Mistral-small | 77.3 | 88.4 | 73.6 | 88.4 | 86.9 | 88.4 | 75.1 | 88.4 | 71.6 | 88.4 | 87.8 | 89.5 | |
| Mistral-large | 79.5 | 88.8 | 75.3 | 88.4 | 85.0 | 88.8 | 75.8 | 88.9 | 72.4 | 88.4 | 83.8 | 87.5 | |
| Claude-3 | 73.5 | 88.4 | 66.4 | 88.4 | 87.6 | 88.9 | 76.4 | 88.4 | 72.6 | 88.5 | 89.3 | 89.3 | |
| Claude-3.5-v1 | 70.5 | 88.4 | 69.6 | 88.4 | 89.4 | 88.4 | 76.1 | 89.5 | 72.8 | 89.7 | 91.0 | 91.5 | |
| Claude-3.5-v2 | 70.5 | 88.4 | 72.5 | 88.4 | 91.7 | 89.2 | 76.1 | 89.2 | 73.3 | 89.4 | 91.9 | 90.8 | |
| Average | 76.8 | 88.6 | 75.1 | 88.4 | 80.9 | 87.4 | 76.6 | 88.6 | 73.4 | 88.7 | 82.9 | 88.0 | |
| Std Dev | 4.8 | 0.3 | 6.0 | 0.2 | 9.4 | 2.3 | 2.9 | 0.5 | 2.8 | 0.7 | 7.0 | 2.6 | |
| WVS | OLMo-2-7B-Base | 74.5 | 81.2 | 70.7 | 80.3 | 84.5 | 84.4 | 78.2 | 81.1 | 78.2 | 81.0 | 80.8 | 83.0 |
| OLMo-2-7B-SFT | 69.7 | 79.1 | 68.1 | 79.2 | 87.2 | 85.6 | 72.9 | 80.3 | 73.1 | 80.8 | 84.9 | 87.0 | |
| OLMo-2-7B-DPO | 66.2 | 78.3 | 70.5 | 79.7 | 89.1 | 82.8 | 70.6 | 80.3 | 70.2 | 80.8 | 82.5 | 84.3 | |
| OLMo-2-7B-Instruct | 75.3 | 80.9 | 74.9 | 80.5 | 86.6 | 87.6 | 58.3 | 78.1 | 60.0 | 76.9 | 86.0 | 88.6 | |
| Llama-3-70B | 59.7 | 76.3 | 61.7 | 79.3 | 81.3 | 82.9 | 67.9 | 79.2 | 62.8 | 79.6 | 78.1 | 79.4 | |
| Llama-3.1-70B | 61.9 | 76.8 | 56.4 | 75.8 | 65.2 | 70.7 | 63.6 | 78.0 | 57.6 | 76.7 | 68.3 | 72.1 | |
| Llama-3.2-1B | 80.3 | 80.3 | 80.3 | 80.3 | – | – | 68.7 | 82.5 | 66.6 | 79.0 | – | – | |
| Llama-3.2-11B | 75.4 | 81.9 | 76.0 | 84.4 | 82.7 | 83.6 | 78.2 | 83.2 | 78.0 | 81.9 | 66.5 | 77.0 | |
| Llama-3.2-90B | 61.6 | 76.8 | 59.1 | 75.4 | 64.6 | 70.2 | 62.1 | 77.7 | 59.5 | 76.5 | 67.7 | 71.5 | |
| Qwen-2.5-72B | 39.6 | 71.7 | 42.4 | 74.3 | 74.0 | 74.3 | 49.1 | 71.8 | 49.4 | 74.7 | 73.4 | 73.8 | |
| Mistral-small | 42.9 | 73.0 | 46.3 | 74.4 | 75.0 | 75.8 | 48.0 | 72.5 | 46.4 | 74.6 | 68.6 | 71.0 | |
| Mistral-large | 48.8 | 74.2 | 44.0 | 76.4 | 72.8 | 74.8 | 54.3 | 74.4 | 51.5 | 76.4 | 76.6 | 78.3 | |
| Claude-3 | 47.3 | 71.4 | 55.7 | 76.3 | 74.1 | 74.0 | 55.3 | 74.8 | 57.3 | 77.1 | 73.2 | 73.3 | |
| Claude-3.5-v1 | 44.3 | 71.2 | 49.6 | 74.6 | 75.7 | 77.7 | 58.7 | 76.7 | 54.8 | 78.1 | 73.3 | 75.5 | |
| Claude-3.5-v2 | 46.5 | 72.1 | 51.0 | 75.5 | 75.2 | 75.2 | 61.0 | 78.2 | 56.8 | 78.4 | 75.6 | 75.6 | |
| Average | 59.6 | 76.3 | 60.4 | 77.8 | 77.7 | 78.5 | 63.1 | 77.9 | 61.5 | 78.2 | 75.4 | 77.9 | |
| Std Dev | 13.8 | 3.9 | 12.4 | 3.0 | 7.7 | 5.8 | 9.5 | 3.4 | 9.9 | 2.3 | 6.4 | 5.8 | |
| Model | Base prompt | SD prompt | |||
| WGM | Llama-3-70B | 73.1 | 84.6 | 67.5 | 86.1 |
| Llama-3.1-70B | 74.0 | 81.8 | 70.4 | 86.1 | |
| Llama-3.2-1B | 83.6 | 83.5 | 83.0 | 83.1 | |
| Llama-3.2-11B | 84.0 | 84.3 | 82.2 | 81.8 | |
| Average | 78.7 | 83.5 | 75.8 | 84.3 | |
| Std Dev | 5.9 | 1.3 | 8.0 | 2.2 | |
| OQA | Llama-3-70B | 75.4 | 88.4 | 72.3 | 88.5 |
| Llama-3.1-70B | 78.9 | 88.4 | 75.7 | 88.4 | |
| Llama-3.2-1B | 88.3 | 85.2 | 87.3 | 86.6 | |
| Llama-3.2-11B | 82.6 | 88.3 | 88.3 | 88.4 | |
| Average | 81.3 | 87.6 | 80.9 | 88.0 | |
| Std Dev | 5.5 | 1.6 | 8.1 | 0.9 | |
| WVS | Llama-3-70B | 54.1 | 74.8 | 51.1 | 72.1 |
| Llama-3.1-70B | 43.5 | 71.3 | 49.1 | 70.8 | |
| Llama-3.2-1B | 83.9 | 82.1 | 85.1 | 84.6 | |
| Llama-3.2-11B | 74.8 | 78.3 | 72.9 | 75.9 | |
| Average | 64.1 | 76.6 | 64.5 | 75.8 | |
| Std Dev | 18.5 | 4.6 | 17.4 | 6.2 | |
Appendix B Prompts (Reproducibility)
B.1 Demographics used
Demographics are shown in Table 11.
| Dataset | Example Questions | Demographics |
| WGM | How much do you, personally, know about science? Do you know a lot, some, not much, or nothing at all? | age, education, employment status, gender, income quintile, living area (city or rural), world region |
| How much do you trust doctors and nurses in your country? | ||
| OQA | How important, if at all, is being a gun owner to your overall identity? | age, education, income, marital status, political ideology, political party, race, US region, religion, sex |
| How much confidence, if any, do you have in elected officials to act in the best interests of the public? | ||
| WVS | Do you agree strongly, agree, disagree, or disagree strongly with the following statement? One of my main goals in life has been to make my parents proud. | age, education, employment status, household size, immigrant status, income, marital status, religion, sex |
| Do you agree, disagree or neither agree nor disagree with the following statement? When jobs are scarce, employers should give priority to people of this country over immigrants. |
B.2 Base prompts
Base prompts are shown in Table 12.
| Prompt type | Prompt template |
| Verbalized | What percentage of people would respond with {answer choices} to the following question? Question: ‘{q}’ Respond with a list of {l} numbers which represent the percentages for {answer choices} in that order. Expected response format: [, , , ,] Response: |
| Log probabilities | Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: {answer} |
| Self-random | Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 1 | Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 2 | How might someone respond to the following question? Question: ‘{q}’ Possible responses are {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 3 | Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 4 | Below you will be asked to answer some questions. Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 5 | Answer the following question. Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
B.3 Sociodemographic prompts
Sociodemographic prompts are shown in Table 13.
| Prompt type | Prompt template |
| Verbalized | What percentage of people whose {att} is {c} would respond with {answer choices} to the following question? Question: ’{q}’ Respond with a list of {l} numbers which represent the percentages for {answer choices} in that order. Expected response format: [, , , ,] Response: |
| Log probabilities | Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: {answer} |
| Self-random | Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 1 | Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 2 | How would a person whose {att} is {c} respond to the following question? Question: ‘{q}’ Possible responses are {answer choices } Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 3 | Person: {c} {att} Question: ‘{q}’ Response options: {answer choices } Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 4 | Below you will be asked to provide a short description of your {att} and then answer some questions. Description: I am {c} Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 5 | Answer the following question as if your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |