### v2

# Improving the Distributional Alignment of LLMs using Supervision

Abstract
The ability to accurately align LLMs with human population groups on subjective questions would have great value. In this work, we show that use of simple supervision can greatly improve language model alignment with diverse population groups more consistently, as measured over three datasets spanning various topics. Beyond evaluating average alignment, we also report how alignment varies across specific groups. Our broad findings provide insights into the distributional alignment of LLMs with diverse population groups. By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a benchmark to stimulate future research. [^1] Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment
1 Introduction
Figure 1: Prior work studies using persona/sociodemographic prompting to align LLMs with humans for subjective questions. In this work, we elicit distributions from LLMs and calibrate them to better align with human response distributions.
As LLMs are increasingly incorporated into real-world systems, it is more and more important to evaluate their alignment with human values and opinions. A growing body of work seeks to align LLMs with particular human population groups, with the goal of simulating their opinions or behaviors, by steering LLM generations towards those groups Aher et al. (2023); Sun et al. (2024); Dillion et al. (2023). One common method to do so is persona or sociodemographic (SD) prompting, which refers to defining a persona with demographic or description variables within a prompt, with the goal of modeling the behavior of that group using an LLM Hu and Collier (2024); Beck et al. (2024).
Recent work in NLP has sought to evaluate SD prompting by studying its performance on large surveys and polls, which contain many subjective questions, along with hundreds of thousands of human responses Santurkar et al. (2023); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). However, evaluation of SD prompting has typically suffered from one or more key shortcomings. First, many studies have assumed a single majority answer to each question for each sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. Prompting for a distribution, rather than a representative response, could potentially also make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2025a); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). However, Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024), or evaluate LLM distributions prompted for specific demographics in a demographic-agnostic manner Beck et al. (2024).
In this work, we evaluate SD prompting in a distributional manner across three survey datasets using a variety of methods to extract distributions from LLMs. We focus on methods that work most broadly, supporting use with open-source, open-model, and API-access LLMs. We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups. We use supervised calibration to mean learning a transformation to a distribution to be more similar to human response distributions, using human responses as supervised labels (described in detail in Section 3.6). We use alignment to mean the degree to which opinions align between humans and LLMs (see Section 3.4).
Concretely, we investigate distributional opinion alignment via the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data? Is this consistent across models, datasets, and distribution elicitation methods? (RQ 3) Can generated distributions be more easily aligned with some SD groups over others? What is the effect of calibration for individual groups? (RQ 4) Does calibration work with fewer supervised training examples? And (RQ 5) What are the effects of post-training (doing fine-tuning and reinforcement learning) on LLM-generated distributions?
We evaluate the alignment of LLM-generated distributions to human opinion across 3 distribution elicitation methods and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We first show that baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1). In contrast, our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across prompt methods, elicitation methods, datasets, and LLMs (RQ2), though it improves for some SD groups and degrades for others (RQ3). Moreover, as few as 1-10 gold examples can be used to calibrate distributions (RQ4). In regard to LLM post-training (RQ5), calibration results in more consistent behavior across various post-training methods, both with and without SD prompting; without calibration, alignment is much more dependent on both the dataset and the post-training method used.
We conclude with a discussion of the methods and models that generate the distributions most aligned with human groups on subjective questions. Caution is warranted given varying success in aligning to different groups as well as recognition that such efforts risk essentializing sociodemographic categories if not applied with care.
2 Related Work
Social alignment of LLMs
LLMs are increasingly used to simulate social human groups in a variety of settings, such as psychology, sociology, and economic studies Aher et al. (2023); Dillion et al. (2023), survey development Rothschild et al. (2024); Jansen et al. (2023), NLP annotation Nasution and Onan (2024); Hashemi et al. (2024); Movva et al. (2024), and multi-agent interactive simulation Hou et al. (2025); Zhou et al. (2024).
Many methods have been proposed to better align LLMs with human values and opinions Ma et al. (2024). For example, persona or sociodemographic (SD) prompting defines a persona with demographic or description variables within a prompt to simulate or model groups Hu and Collier (2024); Beck et al. (2024). Other methods include: using knowledge of a user’s previous opinions to model their future opinions Hwang et al. (2023); Gordon et al. (2022), using psychological scaffolds to improve model rationalization Joshi et al. (2025), self-alignment via the simulation of social scenes Pang et al. (2024), extracting persona vectors to mitigate certain character traits Chen et al. (2025), creating more descriptive value profiles Sorensen et al. (2025), using in-context learning Choenni and Shutova (2025), and fine-tuning LLMs followed by collection of a small amount of human data Krsteski et al. (2025). Evaluation frameworks have also been proposed for assessing alignment and social simulation Kang et al. (2025); Shi et al. (2025); Lutz et al. (2025).
Pitfalls of social alignment
Various limitations have been reported, such as LLMs being unreliable as personalized annotators Dong et al. (2024), unable to use different languages to steer generation towards groups Kwok et al. (2024), biased towards some groups Sun et al. (2025b), lacking stability and steerability under prompt variations Khan et al. (2025), and lacking the reasoning depth to simulate humans in social experiments Gao et al. (2025). Prior work evaluating SD prompting has also reported mixed results in accurately simulating social groups Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024); Kaiser et al. (2025). While some works show that SD prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024); Bisbee et al. (2024).
Other work warns against such methods due to misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2025a); Cheng et al. (2023b); Kitadai et al. (2024). Simply incorporating demographic features into prompts does not always enable LLMs to adopt the desired perspectives of specified demographic groups Sun et al. (2025a). Furthermore, a vast body of work has established biases in language models that may be perpetuated with SD prompting Nadeem et al. (2021); Gallegos et al. (2024); Gupta et al. (2024).
Probabilistic reasoning with LLMs
Prior work in NLP has studied the use of LLMs for eliciting and working with uncertainty and probability distributions. Paruchuri et al. (2024) study whether LLMs can reason about probability distributions, finding that LLMs are able to answer questions about distributions, though performance varies based on the type of distribution. Wang et al. (2025b) develop a model and framework for modeling and evaluating fine-grained probability estimation of propositions conditioned on context, and show that their model captures some amount of human uncertainty. Prior work has also found that several LLMs are capable few-shot regressors, often outperforming traditional supervised methods Vacareanu et al. (2024), and capable fine-tuned regressors Song et al. (2024).
Distribution prediction in the real world is a complex task, and can rely on both uncertainty quantification and nuanced language understanding Wang et al. (2025b); Vedula et al. (2025). This motivates using LLMs for this task, and prior work has developed methods in various domains for doing LLM-based distribution prediction: combining LLM-generated distributions to better predict human annotator distributions Hashemi et al. (2024), developing a quantile regression approach for price prediction Vedula et al. (2025), and mapping LLM token predictions to continuous distributions for probabilistic forecasting Gruver et al. (2023).
Distributional alignment
Recent works have also advocated for a shift to pluralistic alignment due to the ecological fallacy of assuming all members of a group share the same values, beliefs, or opinions Gordon et al. (2022); Sorensen et al. (2024); Feng et al. (2024); Kirk et al. (2024); Xie et al. (2025). Closer to our work, prior works have studied aligning LLM distributions with human response distributions on subjective questions. We draw inspiration from Santurkar et al. (2023), though we evaluate multiple methods of extracting distributions from LLMs. Dominguez-Olmedo et al. (2024) study systematic biases with LLM-generated distributions and how this differs from human response distributions for surveys. Meister et al. (2025) study different distribution elicitation methods for subjective questions. Cao et al. (2025) fine-tune with a divergence-based loss function. Our work introduces the notion of supervised calibration in this space, and studies the effect of alignment methods on individual groups.
3 Approach
3.1 Datasets
We consider three survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA) Santurkar et al. (2023), and the World Values Survey (WVS). OQA and WVS have been used to study alignment in prior work Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). These three datasets cover diverse subjective topics, including global perceptions of science and public health (WGM), US public opinion (OQA), and various moral opinions and values (WVS). We include all WGM ordinal questions and two questions per category from OQA and WVS, yielding 92 questions across WGM (14), OQA (38), and WVS (40) and a total of 4,500 human response distributions over all datasets, questions, and demographics.
Table 13 shows example questions and demographics for each dataset. For each question, model, and elicitation method (Section 3.2), we use an LLM to generate distributions over answer choices, yielding 220,500 SD-specific response distributions and 4,500 SD-agnostic response distributions.
3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To generate distributions via LLM, we apply (with some modifications) methods inspired from model confidence literature. We investigate techniques that work most broadly, supporting use with open-source, open-model, and API-access LLMs. That said, Appendix A.2 reports a smaller scale study with log probabilities. Appendix B shows prompts used for all methods.
Verbalized. This is directly prompting the model to output a probability distribution Geng et al. (2024); Tian et al. (2023). For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over the answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample $n=3$ times and average for variance reduction. If a model ever does not generate a distribution as requested, we simply discard that output. We also renormalize generated distributions that do not sum to 1.
Self-random. We prompt the model to output a single answer choice, sampling $n$ times Xiong et al. (2024), with a temperature of 0.7. We then create a distribution over the $n$ responses. For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice, such as 4 for a lot or 1 for not at all. Note that this method estimates the log probabilities for large $n$ Tian et al. (2023). In our experiments, we sample $n=5$ times, following prior work Xiong et al. (2024).
Paraphrase. We also prompt the model to output a single answer choice but use different paraphrases of the prompt instead of sampling responses for the same prompt Xiong et al. (2024). We use $n=5$ paraphrases for each prompt.
3.3 Prompts
We use two categories of prompts, with three variations. For each, we prompt with the 3 different elicitation methods. See Appendix B for the prompts.
Base. Base prompts exclude any sociodemographic information, asking questions from the original dataset with formatting slightly changed for LLMs. See Appendix Table 14 for full prompts.
Sociodemographic (SD). These use a form similar to “Imagine you are {d}. Question: ”, where d is a demographic value. The exact prompt varies with demographic and elicitation method. See Appendix Table 15 for full prompts.
Prompt variations. We study three prompt variations, standard prompts, few shot prompts (Section 6.2), and chain of thought (CoT) prompts Kojima et al. (2022); Wei et al. (2022) (Section 4.3). See Appendix Table 16 for the variations.
3.4 Metric: Opinion Alignment
We use the following opinion alignment metric from Santurkar et al. (2023) to measure similarity between an elicited distribution ($D_{1}$) and ground truth distribution ($D_{2}$) for the set of questions $Q$:
$A(D_{1},D_{2};Q)=\frac{1}{|Q|}\sum_{q\in Q}{\frac{1-WD(D_{1}(q),D_{2}(q))}{N-1}}$
where $WD$ refers to Wasserstein distance, or earth-mover’s distance, and N1N-1 is a normalization factor. Santurkar et al. (2023) note this metric accounts for the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges in $[0,1]$, but in our results we show the value as a percentage.
3.5 Models
We evaluate 15 models ranging in openness (open-source, open-weight, black box), size, modality, and post-training. Model families include Claude, Llama, Mistral, OLMo-2, and Qwen. Our main results report on only the most powerful model in each family. Others are reported in Appendix A.1.
Model Base prompt Sociodemographic prompt
$P$ $P_{C}$ $S$ $S_{C}$ $V$ $V_{C}$ $P$ $P_{C}$ $S$ $S_{C}$ $V$ $V_{C}$
WGM Claude-3.5-v2 66.2 85.2 59.5 85.1 89.3 87.7 65.4 84.4 61.4 84.2 89.0 89.3
Llama-3.2-90B 68.1 84.6 73.0 86.2 84.8 89.0 70.6 86.0 67.6 86.1 85.0 89.8
Mistral-large 62.4 84.7 72.0 83.9 89.4 88.9 68.4 84.7 63.0 84.9 87.3 88.4
OLMo-2-7B-I 59.6 81.5 67.7 80.4 59.8 85.1 62.3 82.5 64.5 82.7 69.8 84.6
Qwen-2.5-72B 57.7 83.1 53.3 83.2 88.2 87.1 66.0 84.1 63.4 85.2 89.1 89.4
Average 62.8 83.8 65.1 83.8 82.3 87.6 66.5 84.3 64.0 84.6 84.0 88.3
OQA Claude-3.5-v2 70.5 88.8 72.5 90.4 91.7 91.7 76.1 89.9 73.3 89.6 91.9 91.6
Llama-3.2-90B 79.3 89.6 75.3 89.4 86.8 87.9 79.2 90.1 76.1 90.0 83.4 85.9
Mistral-large 79.5 89.9 75.3 88.3 85.0 86.2 75.8 89.2 72.4 89.5 83.8 84.7
OLMo-2-7B-I 72.6 89.0 72.3 88.4 65.4 79.9 72.8 88.4 70.5 88.6 68.5 81.5
Qwen-2.5-72B 73.9 89.7 67.0 88.8 88.4 87.9 74.4 90.0 71.3 89.5 89.2 88.6
Average 75.2 89.4 72.5 89.1 83.5 86.7 75.7 89.5 72.7 89.4 83.4 86.5
WVS Claude-3.5-v2 46.5 80.3 51.0 80.3 75.2 80.3 61.0 80.4 56.8 80.4 75.6 81.7
Llama-3.2-90B 61.6 79.9 59.1 80.3 64.6 80.3 62.1 80.8 59.5 81.5 67.7 82.7
Mistral-large 48.8 80.3 44.0 82.2 72.8 80.3 54.3 80.4 51.5 80.3 76.6 83.8
OLMo-2-7B-I 75.3 80.3 74.9 80.3 86.6 86.5 58.3 79.2 60.0 77.2 86.0 89.8
Qwen-2.5-72B 39.6 82.0 42.4 82.3 74.0 77.9 49.1 82.4 49.4 81.5 73.4 81.7
Average 54.4 80.6 54.3 81.1 74.6 81.1 57.0 80.6 55.4 80.2 75.9 83.9
Average Claude-3.5-v2 61.1 84.8 61.0 85.3 85.4 86.6 67.5 84.9 63.8 84.7 85.5 87.5
Llama-3.2-90B 69.7 84.7 69.1 85.3 78.7 85.7 70.6 85.6 67.7 85.9 78.7 86.1
Mistral-large 63.6 85.0 63.8 84.8 82.4 85.1 66.2 84.8 62.3 84.9 82.6 85.6
OLMo-2-7B-I 69.2 83.6 71.6 83.0 70.6 83.8 64.5 83.4 65.0 82.8 74.8 85.3
Qwen-2.5-72B 57.1 84.9 54.2 84.8 83.5 84.3 63.2 85.5 61.4 85.4 83.9 86.6
Average 64.1 84.6 64.0 84.6 80.1 85.1 66.4 84.8 64.0 84.7 81.1 86.2
Table 1: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase ($P$), self-random ($S$), and verbalized ($V$). Bolded values are significant between each pair (calculated via paired t-test and Bonferroni correction). The best performance for each model is highlighted in yellow. The best averages across models and datasets are highlighted in blue. Results for all LLMs are shown in Appendix Table 7. Adding SD information does not consistently improve alignment, but calibration more consistently improves alignment on average and in most settings.
3.6 Calibration
We apply supervised regression to transform LLM-generated distributions to more closely align them with the ground truth human distributions. For example, for the question, How much do you trust doctors and nurses? A lot, some, not much, or not at all?, the LLM might generate a distribution such as $[0.6,0.2,0.2,0]$, while the ground truth human response distribution based on the survey question might be $[0.25,0.5,0.1,0.15]$. We seek to learn a transformation from the LLM-generated distributions to the human distributions that might be applied to all LLM-generated distributions. More formally, for a regression model $R$ and LLM distribution $\mathbf{D}$ parameterized by the values for each answer choice $D_{a},D_{b},D_{c},...,D_{k}]$ (where $k$ is the number of answer choices), we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice $G_{a},G_{b},G_{c},...,G_{k}]$. We then re-normalize the transformed distribution.
Training examples. The examples for the regression model are the distributions split by answer choices. For example, the LLM and human distributions given above, $[0.6,0.2,0.2,0]$ and $[0.25,0.5,0.1,0.15]$, would be split into supervised pairs of $(0.6,0.25)$, $(0.2,0.5)$, $(0.2,0.1)$, and $(0,0.15)$. We thus train $R$ on $(X,y)$ pairs of $(D_{a},G_{a}),(D_{b},G_{b}),$... to learn transformed values $D_{a^{\prime}},D_{b^{\prime}},D_{c^{\prime}},...,D_{k^{\prime}}]$ on held-out test questions.
Regression model. We learn a regression model for each dataset-LLM-elicitation-prompt setting (270 regression models for 3 datasets, 15 models, 3 elicitation methods, and 2 prompt categories). Hyperparameter tuning and model selection are used to choose between linear regression (including lasso and ridge) and random forest regression for each regressor. We use the scikit-learn Pedregosa et al. (2011) implementation for each regressor.
Training. The total number of examples for each regression model is $C$ x $Q$ x $A$, where $C$ is the class within each sociodemographic group (e.g., female and male for sex), $Q$ is the question (e.g., how much do you trust doctors?), and $A$ is the number of answer choices per question. These values are different for each setting (see Table 13 for the sizes of $Q$ and $C$ per dataset). We were unable to extract distributions for all questions from some of the smaller LLMs (see Table 17), so the number of training examples differs for them. We split the LLM distributions into train, development, and test sets (80%-10%-10%).
For minimal supervision (Section 6), we train on a random sample of $n$ examples ($n$ = [1, 5, 10, 50, 100, 200, 500, 1000, 1200], where each example is the full distribution not split by answer choices) from the training set. We repeat this procedure 10 times for different random splits and average.
Evaluation. We optimize the regression models with Mean Squared Error (MSE). Since we split distributions during training, we normalize output values so that they sum to 1 before calculating opinion alignment with the ground truth distribution. For example, if the output of $R$ for the distribution $[0.6,0.2,0.2,0]$ is $[0.3,0.7,0.1,0.1]$, we would normalize these values to $[0.25,0.59,0.08,0.08]$.
4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts and calibrating LLM-generated distributions with human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The alignment values for all models are shown in Appendix Table 7 for all the results in this section. All significance results are calculated via paired t-test with Bonferroni correction.
4.1 Does SD prompting improve alignment?
For RQ1, we compare SD vs. base prompts to study the effect of adding SD information on opinion alignment. Note that while prior work studies this question only for majority-voted responses Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024), we investigate the effect of adding SD information when comparing distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets.
For base and SD prompts, LLM distributions are generally most aligned with verbalized elicitation (Table 1), with exceptions (e.g., on the OQA dataset). As shown in the table, prompting with SD information often leads to comparable or even lower opinion alignment than prompting without any SD information, although this is model, dataset, and elicitation method dependent.
4.2 Can we calibrate LLM distributions?
We have seen that prompting LLMs with SD information does not consistently improve opinion alignment with human responses. We next investigate for RQ2 whether calibrating these distributions can more accurately and consistently align them with human response distributions.
The results after regression, compared to pre-regression results, are shown in Table 1. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, improving by an average of 16.3%, where we average the delta between uncalibrated and calibrated opinion alignment across dataset-LLM-elicitation method settings. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment in these settings decreases by an average of 4.1%. These results suggest that LLM-generated distributions are somewhat uncalibrated, and a simple supervised regression can improve alignment.
Figure 2: Standard deviation vs. opinion alignment. Each point represents the average alignment for each dataset, LLM, and elicitation method. For visual clarity, we omit 43/290 uncalibrated points having opinion alignment below 60. Calibration tends to both increase opinion alignment and decrease standard deviation. It also decreases variance between settings.
Model Base prompt Sociodemographic prompt
$P$ $P_{cot}$ $P_{C}$ $S$ $S_{cot}$ $S_{C}$ $V$ $V_{cot}$ $V_{C}$ $P$ $P_{cot}$ $P_{C}$ $S$ $S_{cot}$ $S_{C}$ $V$ $V_{cot}$ $V_{C}$
WGM Claude 66.2 65.6 85.2 59.5 60.7 85.1 89.3 88.8 87.7 65.4 61.7 84.4 61.4 58.6 84.2 89.0 79.7 89.3
OLMo 59.6 62.9 81.5 67.7 62.7 80.4 59.8 86.1 85.1 62.3 64.0 82.5 64.5 59.8 82.7 69.8 81.8 84.6
OQA Claude 70.5 76.4 88.8 72.5 74.3 90.4 91.7 89.9 91.7 76.1 68.0 89.9 73.3 66.1 89.6 91.9 80.3 91.6
OLMo 72.6 79.3 89.0 72.3 77.0 88.4 65.4 85.2 79.9 72.8 77.1 88.4 70.5 76.3 88.6 68.5 82.9 81.5
WVS Claude 46.5 55.5 80.3 51.0 52.8 80.3 75.2 80.3 80.3 61.0 61.4 80.4 56.8 59.8 80.4 75.6 79.7 81.7
OLMo 75.3 76.9 80.3 74.9 73.5 80.3 86.6 86.5 86.5 58.3 72.7 79.2 60.0 69.4 77.2 86.0 85.7 89.8
Table 2: Opinion alignment for original, CoT prompted, and calibrated distributions for base and SD prompts. Each set of three columns compares the elicitation methods: paraphrase ($P$), self-random ($S$), and verbalized ($V$). Bolded values are significantly better than others in each triple (calculated via Anova followed by paired t-tests and Bonferroni correction). The highlighted value is the highest alignment for each model. Claude refers to Claude-3.5-v2 and OLMo refers to OLMo-2-7B-Instruct. CoT significantly improves alignment for verbalized elicitation for OLMo. For Claude, verbalized alignment is not significantly different or drops with CoT.
After calibration, opinion alignment in each setting also has much lower variance. Calibrated distributions have a lower standard deviation in 87.2% of dataset-LLM-elicitation settings, and are 1.62 times lower on average. That lower standard deviation leads to higher opinion alignment provides further evidence that LLMs may exaggerate differences in opinion distributions between demographic groups Cheng et al. (2023a, b), and that calibration could help mitigate this.
We also find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. Standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. Although there are some settings where uncalibrated distributions have a lower standard deviation than their calibrated counterparts, we emphasize that the variance between settings is much lower after calibration. While prior work Beck et al. (2024); Hu and Collier (2024) and our own results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly. This means that with calibration, the choice of any particular model becomes less important.
We do not find distinct patterns with model size within model family (e.g., Mistral small vs. Mistral large). Indeed, smaller models are sometimes more aligned than larger models, though across prompting methods, elicitation methods, and datasets, larger models are more aligned more consistently (see Appendix Table 7). However, we do find that chain of thought prompting (Section 4.3) appears to increase alignment for smaller models, but not for larger models. The smaller Llama models in particular often refused to answer questions (i.e., questions would hit safety restrictions for attempting to role-play as a particular identity group), or would not follow instructions in formatting output distributions. We were thus unable to extract distributions for all questions, leading to higher variance in alignment. See Appendix Table 17 for percent distributions extracted in each setting.
Demographic Claude-3.5-v2 Llama-3.2-90B Mistral-large OLMo-2-7B-I Qwen-2.5-72B Average
$V$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $V_{C}$
World region (WGM) Aus/NZ 82.3 79.7 94.7 83.5 88.2 86.6 73.5 79.1 89.1 84.2 85.6 82.6
Central Africa 85.5 87.7 71.2 81.2 78.1 78.6 38.0 79.0 75.6 79.2 69.7 81.1
Cent. America & Mex. 89.2 90.0 74.8 85.1 81.7 82.8 46.4 84.7 78.5 81.8 74.1 84.9
Central Asia 82.5 80.2 88.9 82.8 87.7 85.3 75.5 79.9 88.8 84.4 84.7 82.5
East Asia 88.4 85.7 86.8 85.7 88.7 88.6 61.6 82.0 86.4 82.8 82.4 85.0
Eastern Africa 88.2 88.4 80.0 87.7 88.0 87.3 63.3 87.7 86.0 86.2 81.1 87.5
Eastern Europe 93.0 88.4 85.4 88.1 89.6 89.4 54.6 81.9 89.9 88.5 82.5 87.3
Middle East 91.1 87.9 87.7 92.9 92.5 91.7 59.0 85.1 91.4 88.8 84.3 89.3
North Africa 90.0 87.5 81.6 87.8 86.9 87.2 53.0 81.4 85.1 84.1 79.3 85.6
Northern America 83.2 80.9 91.7 85.0 88.2 87.0 68.4 82.4 92.4 88.8 84.8 84.8
Northern Europe 82.8 80.6 94.7 84.8 89.6 87.6 72.4 81.0 91.3 86.5 86.2 84.1
South America 88.9 89.5 74.0 84.4 80.9 82.0 45.1 82.5 77.7 81.0 73.3 83.9
South Asia 84.2 84.0 86.7 84.8 88.9 87.3 70.5 83.7 87.0 83.3 83.5 84.6
Southeast Asia 82.1 82.5 79.1 80.2 86.3 84.6 70.6 81.9 84.2 81.7 80.5 82.2
Southern Africa 86.4 90.1 74.2 84.4 82.4 82.8 49.1 88.8 79.7 83.3 74.4 85.9
Southern Europe 93.2 89.1 84.5 88.7 90.0 89.7 53.4 81.7 88.5 88.2 81.9 87.5
Western Africa 90.7 90.4 81.0 91.6 88.1 88.2 57.7 90.3 85.7 86.4 80.6 89.4
Western Europe 83.2 80.6 93.0 86.4 90.3 88.5 71.6 84.2 93.5 88.7 86.3 85.7
All Demographics (WGM) 89.3 87.7 84.8 89.0 89.4 88.9 59.8 85.1 88.2 87.1 82.3 87.5
P.I.(OQA) Very conservative 85.1 84.8 77.4 79.0 76.1 77.9 63.8 77.5 81.7 81.0 76.8 80.0
Conservative 89.0 89.0 82.4 83.7 81.3 82.9 64.4 79.2 85.7 85.1 80.6 84.0
Moderate 92.3 92.3 87.9 88.9 85.7 87.0 65.0 79.5 88.5 88.0 83.9 87.1
Liberal 91.6 92.1 89.5 89.3 85.5 85.9 63.1 77.8 88.6 88.9 83.7 86.8
Very liberal 87.8 88.3 86.5 85.9 83.3 83.3 61.1 76.1 84.8 85.0 80.7 83.7
All Demographics (OQA) 91.7 91.7 86.8 87.9 85.0 86.2 65.4 79.9 88.4 87.9 83.5 86.7
I.(WVS) High 77.1 82.2 66.0 82.2 74.3 82.2 87.6 89.9 75.9 79.8 76.2 83.3
Middle 75.2 80.2 64.7 80.2 72.7 80.2 86.8 86.6 73.6 77.8 74.6 81.0
Low 77.4 82.4 66.8 82.4 75.0 82.4 87.8 88.5 76.1 80.0 76.6 83.1
All Demographics (WVS) 75.2 80.3 64.6 80.3 72.8 80.3 86.6 86.5 73.9 77.9 74.6 81.1
Table 3: Opinion alignment before ($V$) and after ($V_{C}$) calibration for three demographic categories (one from each dataset) using base-prompted, verbalized elicitation. “P.I” is political ideology and “I.” is income. Significantly higher values between each pair of columns (calculated via a paired t-test and Bonferroni correction) are bolded. The two “Average” columns on the right are averages across models, and the “All Demographics” rows are averages across the total set of SDs per dataset. See Appendix A.3 for all SDs. Calibrated distributions are more aligned with some SDs over others. Some models are better aligned with some datasets, e.g., Claude with OQA SDs and OLMo with WVS SDs.
4.3 The effect of chain of thought prompting
We compare CoT prompting with standard prompting + calibration for two models, the model that is most aligned on average (Claude 3.5 v2) and a small model (OLMo-2-7B Instruct). Our results are shown in Table 2. We find that for Claude, CoT does not consistently improve alignment over standard prompting, though it does improve in some settings. For base prompting with verbalized elicitation, alignment with CoT is not significantly different than with calibration for 2/3 datasets. For SD prompting, calibration creates significantly more aligned distributions for all datasets. For OLMo, we find that CoT improves alignment much more consistently, often leading to distributions that are just as aligned as, or better aligned than, calibrated distributions for the verbalized setting. Calibration is consistently more aligned than CoT for distributions elicited via paraphrase or self-random.
Upon qualitative inspection, we found that SD + CoT prompting Claude led to an “overthinking” effect Pu et al. (2025), where the model sometimes generated a distribution that exaggerated a particular difference with the ground truth. For example, the knowledge that “Men generally report higher confidence in scientific knowledge than women due to societal factors” to generate a distribution for the question “How much do you know about science?” led to the model producing a distribution that was more skewed than the ground truth. We hypothesize that this led to lower alignment for Claude (and would be similar for other larger, more capable reasoning models), but not for smaller models like OLMo 7B. Future work might study this further.
5 How does alignment vary across SDs?
So far, we have studied alignment with SD groups aggregated across groups, in order to compare different models and distribution elicitation methods. However, a particular model or method might be more aligned with some sociodemographics over others. With RQ3, we seek to understand the differences at the more granular demographic level.
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it leads to the most aligned distributions with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 3. We show results for the same 5 LLMs used in Section 4. See Appendix A.3 for alignment for other demographics.
5.1 The effect of calibration
First, we compare opinion alignment between base and calibrated distributions. We use the regression models trained on all SD groups and study how such aggregate regression models might calibrate individual sociodemographics. We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more absolutely aligned than their uncalibrated counterparts for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. This includes all demographic categories, including those in Table 3 and Appendix A.3. The delta improvement across demographics is most significant for OLMo on the WGM and OQA datasets. As expected, calibration improves alignment for some demographics and reduces alignment for others. We find that it tends to improve alignment of SD groups that were less aligned with base-prompted distributions, such as Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment decreases for very high aligned demographic groups, such as Aus/NZ and Northern Europe.
Figure 3: Mean Squared Error (MSE) of regression models on various training data sizes, using SD prompted and verbally elicited distributions. Plots are shown for each dataset: (a) WGM, (b) OQA, and (c) WVS. Although model and dataset dependent, MSE converges between 1 and 10 examples.
Model Base prompt Sociodemographic prompt
$P$ $P_{F}$ $P_{C}$ $S$ $S_{F}$ $S_{C}$ $V$ $V_{F}$ $V_{C}$ $P$ $P_{F}$ $P_{C}$ $S$ $S_{F}$ $S_{C}$ $V$ $V_{F}$ $V_{C}$
WGM Claude 66.2 63.5 86.4 59.5 61.7 83.7 89.3 89.0 89.2 65.4 59.8 84.1 61.4 58.1 84.2 89.0 78.9 89.0
Llama 68.1 72.5 82.8 73.0 66.5 86.4 84.8 86.0 89.2 70.6 66.7 86.0 67.6 64.9 85.6 85.0 78.7 89.6
Mistral 62.4 59.5 84.5 72.0 55.4 83.7 89.4 87.2 89.6 68.4 58.4 85.8 63.0 55.3 84.9 87.3 78.1 88.5
OLMo 59.6 70.5 81.1 67.7 70.9 81.0 59.8 81.1 83.7 62.3 64.2 82.2 64.5 66.0 81.6 69.8 79.3 80.6
Qwen 57.7 60.6 82.8 53.3 59.0 82.1 88.2 88.3 87.3 66.0 58.5 85.2 63.4 55.9 85.4 89.1 80.4 89.4
Average 62.8 65.3 83.5 65.1 62.7 83.4 82.3 86.3 87.8 66.5 61.5 84.7 64.0 60.0 84.3 84.0 79.1 87.4
OQA Claude 70.5 77.9 88.4 72.5 75.7 89.6 91.7 91.5 91.5 76.1 68.1 89.7 73.3 66.9 89.7 91.9 79.7 92.3
Llama 79.3 77.6 89.7 75.3 76.6 89.6 86.8 88.8 88.6 79.2 75.0 90.0 76.1 75.5 89.8 83.4 80.5 87.1
Mistral 79.5 77.4 89.6 75.3 74.0 89.0 85.0 89.7 86.9 75.8 70.5 89.8 72.4 69.9 87.5 83.8 80.7 87.8
OLMo 72.6 77.0 88.9 72.3 74.6 85.9 65.4 81.2 77.4 72.8 72.3 88.6 70.5 74.4 88.4 68.5 78.8 86.6
Qwen 73.9 75.9 90.1 67.0 74.4 89.3 88.4 88.9 87.2 74.4 71.4 90.2 71.3 69.3 89.3 89.2 81.0 87.8
Average 75.2 77.2 89.3 72.5 75.1 88.7 83.5 88.0 86.3 75.7 71.5 89.7 72.7 71.2 88.9 83.4 80.1 88.3
WVS Claude 46.5 62.4 81.7 51.0 62.1 80.8 75.2 82.1 81.5 61.0 64.7 80.4 56.8 64.0 80.2 75.6 81.9 80.4
Llama 61.6 70.1 79.4 59.1 68.9 80.9 64.6 79.7 83.5 62.1 70.3 80.0 59.5 72.2 80.9 67.7 78.8 86.2
Mistral 48.8 63.9 81.4 44.0 61.0 80.6 72.8 81.0 81.1 54.3 65.0 82.4 51.5 62.7 81.4 76.6 78.7 82.5
OLMo 75.3 72.9 80.0 74.9 74.2 80.2 86.6 84.2 87.8 58.3 68.8 78.8 60.0 72.3 77.5 86.0 78.0 89.7
Qwen 39.6 58.3 82.1 42.4 56.8 81.3 74.0 78.4 81.0 49.1 59.0 81.5 49.4 56.8 81.2 73.4 78.3 76.6
Average 54.4 65.5 80.9 54.3 64.6 80.8 74.6 81.1 83.0 57.0 65.6 80.6 55.4 65.6 80.2 75.9 79.1 83.1
Table 4: Opinion alignment for original, few shot promptedF, and calibratedC distributions for base and SD prompts. Each set of three columns compares the three elicitation methods: paraphrase (‘P’), self-random (‘S’), and verbalized (‘V’). Bolded values are significant between each triple (calculated via Anova followed by paired t-tests and Bonferroni correction). The value highlighted in yellow is the absolute best performing overall for each model. The values highlighted in blue are the best averages across models. Each model refers to the largest model we study within the model family. Few shot prompting does not consistently improve alignment across datasets, models, and elicitation methods.
5.2 Alignment variation across LLMs
We now compare opinion alignment between different LLMs. For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East. Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data. For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. This is discussed further in Section 7.
In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others. All models have lowest alignment with middle income populations. This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income. In other words, income might be less important in determining survey answers for those with middle income.
6 Does supervision work with less data?
Our results presented thus far used 80% of each dataset for training our supervised regression models to calibrate the LLM-generated distributions. We next investigate whether smaller amounts of supervision can suffice for opinion alignment (RQ4).
6.1 Amount of supervised data required
As discussed in Section 3.6, we evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset, and average over 10 random samples for each size. We find that as few as 1-10 examples (with $\approx$4 answer choices per example) can suffice to closely match the minimal MSE for any particular dataset-LLM-elicitation setting. Figure 3, plotting Mean Squared Error (MSE) over training data size, shows that MSE usually converges between 1 and 10 examples, though this is model and dataset dependent.
6.2 Comparison with few-shot prompting
Given that the regression models appear to converge with so few examples, one might ask whether the regression is actually learning anything useful, or if we could achieve a similar improvement with few shot prompting instead. To investigate this, we 5-shot prompt the LLMs with the same examples given to the regression model in the 5 example calibration setting. We use a 5-shot prompt since this lies in the middle of how many examples were needed for convergence for the regression models. Whereas we averaged over 10 random samples in those experiments, here we use only 3 of the random samples because we found little variation. We average alignment over the 3 runs.
Table 4 compares few shot prompted ($F$) distributions to original and calibrated ($C$) distributions for base and SD prompts. While 5-shot prompted distributions are not as aligned as calibrated distributions in most settings, they are significantly more aligned in a few cases. However, calibration leads to distributions that are more aligned more consistently, across models, prompting methods, and distribution elicitation methods.
6.3 The effect on individual SD groups
Although we see convergence with fewer examples, does using a small set of random examples affect alignment of individual demographics? We calibrate with a set of 5 examples (in the middle of the 1-10 range described above for MSE convergence). Although degradation on average is close to zero, individual demographics are affected differently. Table 5 shows the five SDs most affected and the degradation amount for both base and SD-prompted verbalized distributions. We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant. Using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are historically less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
7 Does post-training impact alignment?
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + supervised finetuning (SFT), base + SFT + direct preference optimization (DPO), and the instruct model (base + SFT + DPO + reinforcement learning with verifiable rewards (RLVR)). The SFT step involves finetuning on instruction datasets and scaled synthetic data, the DPO step involves preference finetuning with DPO, and the RLVR step involves finetuning targeted at domains where prompts with verifiable rewards can be constructed, such as math OLMo et al. (2025). We use these four models out-of-the-box via Huggingface and study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it is typically most aligned.
Figure 4 plots opinion alignment for all OLMo models. Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets. OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to have less effect in this setting.
We suspect the increase in alignment on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS. The drop in alignment for uncalibrated distributions on WGM and OQA might indicate that OLMo’s post-training data is less aligned with some of the populations of these datasets. Overall, calibration leads to more consistent alignment, with and without SD prompting, regardless of post-training or dataset. This is in agreement with the rest of our results, where using calibration brings more consistency across settings.
Category Demographic $V_{F}$ $V_{5}$ $\Delta$
Base Region Southeast Asia 81.0 79.4 1.6
Education Less than HS 83.1 81.6 1.5
Race Black 86.6 85.1 1.5
Race Hispanic 87.2 85.9 1.4
Pol. Party Democrat 84.7 83.3 1.4
SD Region Northern Europe 87.1 84.6 2.4
Region Western Europe 87.1 84.9 2.3
Region Aus/NZ 84.3 82.1 2.2
Education Tertiary 88.6 86.5 2.1
Employment Unemployed 86.9 84.9 2.0
Table 5: Demographics with largest opinion alignment degradation from calibration models trained on the full dataset ($V_{F}$) vs. only five examples ($V_{5}$). $\Delta$ values shown differences. Distributions are verbally elicited, with base-prompted shown on top and SD-prompted on bottom. Fewer training examples does not lead to any statistically significant degradation. SD prompting changes the groups with most absolute degradation.
Figure 4: Opinion alignment for OLMo-2-7B models with different post-training methods using the verbalized distribution elicitation method. Calibration results in more consistent alignment across all post-training methods, both with and without SD prompting. Without calibration, alignment is much more dataset and post-training method dependent.
8 Discussion
Limitations of demographic modeling. Modeling groups wrt. their demographics assuredly provides only a narrow, incomplete view of people, and demographics alone do not determine our opinions. That said, such models may usefully approximate distributions of group opinions, provided care is exercised. We note that our methods for improving alignment should not be viewed as a way to replace human respondents in surveys, but rather a step towards more accurately representing groups.
The effect of SD prompting. We showed that simply prompting with SD information does not necessarily lead to more aligned distributions. Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al. (2023); Do et al. (2025). One could make personalized predictions for a sample of a group’s individuals members Gordon et al. (2022); Hwang et al. (2023) rather than a single distributive prediction for the group. Multimodal prompts might also add SD information implicitly, though this might come with additional biases.
LLM distributions are uncalibrated. We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated; simple regression allows us to scale distributions to be more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others. We also find that calibration reduces variance, both across examples and across settings. Since calibrated distributions generally have higher performance, this suggests LLMs exaggerate differences between sociodemographic groups, which calibration might help mitigate.
Calibration on individual SD groups. We do not find many significant differences between calibrated/uncalibrated distributions at the individual SD group level. However, we hypothesize that calibrating responses from each individual SD would likely improve its alignment, though we leave this to future work. However, efficacy may vary by demographic; some demographic traits are more relevant to the response distributions of certain questions (e.g., urban vs. rural areas might affect gun rights questions more than marital status).
The choice of LLM and distribution elicitation method matters less with calibration. We find that distributional opinion alignment is dataset, LLM, and distribution elicitation method dependent, consistent with prior work evaluating majority responses Beck et al. (2024). However, verbally-elicited distributions are most aligned with human distributions in the most settings. We also find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is also dataset-dependent. More importantly however, we find that calibration reduces variance in alignment across models and methods; i.e., the choice of model and elicitation method matters less after calibration.
9 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques, as well as increase consistency between settings. Our benchmark also serves to help enable and stimulate future research.
10 Limitations
Prompting methods. We do not experiment with varied temperatures for our generations. For elicitation methods, though a larger $n$ might improve generated distributions, our choice of $n$ is reasonable given inference costs and following prior work Xiong et al. (2024). We use $n=3$ for verbalized elicitation (as opposed to $n=5$) since we found little variance between runs.
Logit-based elicitation methods. We focus on methods that broadly work with open-source, open-model, and API-access LLMs. That said, we report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.2). We found that log probabilities with calibration yielded the most aligned distributions in 10/24 settings considered. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied. We only studied a subset of the demographics available in the datasets. Future work might study additional and/or intersectional demographics Crenshaw (1991).
Regression. Our regression predicts each answer choice individually and then normalizes the predicted answers. We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better. We also optimize regression for MSE instead of the alignment metric. Future work might also ensemble multiple elicitation methods.
Ethical Considerations
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for creating LLMs that more accurately portray and reason about population groups, this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
We note that by prompting LLMs with SD information, and evaluating their distributions against humans with those SDs, we are essentializing the SD to the identities of the survey respondents and their opinions Wang et al. (2025a). Although SDs are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). By seeking to infer distributions of human responses rather than majority responses for each SD group, we seek to align LLMs with groups in a distributionally pluralistic manner, rather than assume all members of the group would answer similarly.
Acknowledgments
We thank Sooyong Lee for early contributions that did not make it into the ultimate paper. This research was supported in part by Cisco, Good Systems [^2] https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), and a grant from Open Philanthropy. We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix A Additional Results
A.1 Alignment results for all models
Results are in Table 7.
A.2 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 6.
Model Base prompt SD prompt
$L$ $_{C} LL$ $L_{C}$
WGM Llama-3-70B 73.1 86.2 67.5 86.9
Llama-3.1-70B 74.0 83.2 70.4 88.1
Llama-3.2-1B 83.6 88.6 83.0 85.1
Llama-3.2-11B 84.0 85.0 82.2 84.4
Average 78.7 85.7 75.8 86.1
Std Dev 5.9 2.3 8.0 1.7
OQA Llama-3-70B 75.4 85.2 72.3 88.1
Llama-3.1-70B 78.9 87.3 75.7 87.4
Llama-3.2-1B 88.3 88.4 87.3 86.5
Llama-3.2-11B 82.6 86.3 88.3 89.6
Average 81.3 86.8 80.9 87.9
Std Dev 5.5 1.4 8.1 1.3
WVS Llama-3-70B 54.1 82.8 51.1 83.5
Llama-3.1-70B 43.5 83.3 49.1 83.0
Llama-3.2-1B 83.9 86.4 85.1 84.8
Llama-3.2-11B 74.8 81.1 72.9 85.5
Average 64.1 83.4 64.5 84.2
Std Dev 18.6 2.2 17.4 1.1
Table 6: Opinion alignment before and after calibration for each dataset and LLM, using log probability distributions. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for log probabilities (‘L’). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
Model Base prompt Sociodemographic prompt
$P$ $P_{C}$ $S$ $S_{C}$ $V$ $V_{C}$ $P$ $P_{C}$ $S$ $S_{C}$ $V$ $V_{C}$
WGM OLMo-2-7B-Base 65.6 80.1 79.2 81.8 82.2 83.9 74.8 81.9 71.0 81.8 74.9 82.9
OLMo-2-7B-SFT 80.6 81.8 70.3 81.8 68.6 86.3 75.6 82.6 70.2 82.1 65.4 85.0
OLMo-2-7B-DPO 71.1 77.4 63.2 82.1 63.9 86.2 72.2 83.0 68.7 82.4 63.8 84.9
OLMo-2-7B-Instruct 59.6 81.5 67.7 80.4 59.8 85.1 62.3 82.5 64.5 82.7 69.8 84.6
Llama-3-70B 64.9 85.3 73.7 83.8 84.5 88.8 66.5 86.0 61.7 84.7 84.4 88.3
Llama-3.1-70B 73.2 86.7 70.3 86.5 80.6 86.0 68.5 85.9 66.1 85.9 85.2 89.4
Llama-3.2-1B 81.8 81.8 81.8 81.8 61.9 82.5 46.7 81.2 61.4 86.4
Llama-3.2-11B 73.7 87.7 73.2 83.5 73.6 84.5 71.8 85.3 72.0 84.7 71.7 79.2
Llama-3.2-90B 68.1 84.6 73.0 86.2 84.8 89.0 70.6 86.0 67.6 86.1 85.0 89.8
Qwen-2.5-72B 57.7 83.1 53.3 83.2 88.2 87.1 66.0 84.1 63.4 85.2 89.1 89.4
Mistral-small 61.6 83.9 61.6 84.6 88.9 88.8 67.5 86.6 64.4 85.8 89.0 89.1
Mistral-large 62.4 84.7 72.0 83.9 89.4 88.9 68.4 84.7 63.0 84.9 87.3 88.4
Claude-3 57.8 81.1 50.0 81.8 84.8 86.7 64.9 80.6 52.9 81.7 88.4 88.1
Claude-3.5-v1 64.5 82.5 62.0 84.3 86.1 87.4 65.8 84.6 62.3 84.5 86.3 88.3
Claude-3.5-v2 66.2 85.2 59.5 85.1 89.3 87.7 65.4 84.4 61.4 84.2 89.0 89.3
Average 67.3 83.2 67.4 83.4 80.3 86.9 68.1 84.0 63.7 83.9 79.4 86.9
Std Dev 7.5 2.7 9.0 1.8 9.9 1.7 4.1 1.8 6.7 1.7 10.3 3.0
OQA OLMo-2-7B-Base 85.0 89.2 82.6 88.4 82.1 81.8 80.8 88.3 77.2 88.1 83.8 87.6
OLMo-2-7B-SFT 78.8 88.0 79.4 86.8 68.7 81.6 80.7 87.9 77.8 87.2 72.6 83.0
OLMo-2-7B-DPO 79.3 88.6 81.5 86.7 69.3 82.8 82.1 88.4 78.7 87.7 72.3 82.7
OLMo-2-7B-Instruct 72.6 89.0 72.3 88.4 65.4 79.9 72.8 88.4 70.5 88.6 68.5 81.5
Llama-3-70B 72.2 87.5 76.8 89.1 81.8 85.0 74.4 88.9 70.1 88.9 79.5 83.7
Llama-3.1-70B 73.5 88.8 70.5 87.7 83.8 86.4 76.4 89.5 71.7 89.4 83.6 86.6
Llama-3.2-1B 82.3 88.4 88.4 88.4 72.6 89.0 70.2 88.6 83.6 84.9
Llama-3.2-11B 84.3 89.4 75.5 86.4 65.8 77.2 75.7 87.2 74.8 87.0 83.0 86.9
Llama-3.2-90B 79.3 89.6 75.3 89.4 86.8 87.9 79.2 90.1 76.1 90.0 83.4 85.9
Qwen-2.5-72B 73.9 89.7 67.0 88.8 88.4 87.9 74.4 90.0 71.3 89.5 89.2 88.6
Mistral-small 77.3 91.6 73.6 90.5 86.9 87.7 75.1 90.4 71.6 90.3 87.8 89.0
Mistral-large 79.5 89.9 75.3 88.3 85.0 86.2 75.8 89.2 72.4 89.5 83.8 84.7
Claude-3 73.5 89.9 66.4 90.3 87.6 86.5 76.4 90.6 72.6 90.3 89.3 88.6
Claude-3.5-v1 70.5 86.9 69.6 88.6 89.4 89.0 76.1 90.3 72.8 89.8 91.0 90.8
Claude-3.5-v2 70.5 88.8 72.5 90.4 91.7 91.7 76.1 89.9 73.3 89.6 91.9 91.6
Average 76.8 89.0 75.1 88.6 80.9 85.1 76.6 89.2 73.4 89.0 82.9 86.4
Std Dev 4.8 1.1 6.0 1.3 9.4 4.0 2.9 1.0 2.8 1.1 7.0 3.0
WVS OLMo-2-7B-Base 74.5 80.3 70.7 80.1 84.5 80.3 78.2 81.4 78.2 81.5 80.8 81.9
OLMo-2-7B-SFT 69.7 79.2 68.1 79.1 87.2 87.2 72.9 80.6 73.1 80.7 84.9 89.0
OLMo-2-7B-DPO 66.2 80.3 70.5 79.8 89.1 83.9 70.6 80.9 70.2 80.7 82.5 85.6
OLMo-2-7B-Instruct 75.3 80.3 74.9 80.3 86.6 86.5 58.3 79.2 60.0 77.2 86.0 89.8
Llama-3-70B 59.7 80.3 61.7 80.6 81.3 83.0 67.9 80.8 62.8 80.5 78.1 82.1
Llama-3.1-70B 61.9 81.5 56.4 82.1 65.2 80.3 63.6 80.8 57.6 81.3 68.3 82.8
Llama-3.2-1B 80.3 80.3 80.3 80.3 68.7 82.8 66.6 78.9
Llama-3.2-11B 75.4 82.0 76.0 84.8 82.7 84.0 78.2 82.2 78.0 81.2 66.5 87.0
Llama-3.2-90B 61.6 79.9 59.1 80.3 64.6 80.3 62.1 80.8 59.5 81.5 67.7 82.7
Qwen-2.5-72B 39.6 82.0 42.4 82.3 74.0 77.9 49.1 82.4 49.4 81.5 73.4 81.7
Mistral-small 42.9 80.3 46.3 80.3 75.0 80.3 48.0 80.4 46.4 80.4 68.6 81.7
Mistral-large 48.8 80.3 44.0 82.2 72.8 80.3 54.3 80.4 51.5 80.3 76.6 83.8
Claude-3 47.3 80.3 55.7 80.3 74.1 80.3 55.3 80.2 57.3 79.8 73.2 81.7
Claude-3.5-v1 44.3 80.3 49.6 80.3 75.7 80.3 58.7 80.3 54.8 80.3 73.3 81.6
Claude-3.5-v2 46.5 80.3 51.0 80.3 75.2 80.3 61.0 80.4 56.8 80.4 75.6 81.7
Average 59.6 80.5 60.4 80.9 77.7 81.8 63.1 80.9 61.5 80.4 75.4 83.8
Std Dev 13.8 0.7 12.4 1.4 7.8 2.7 9.5 0.9 9.9 1.1 6.4 2.9
Table 7: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase (‘P’), self-random (‘S’), and verbalized (‘V’). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section. We note that for some verbalized elicitation settings, Llama-3.2-1B was unable to follow instructions and either did not return any distributions in the format we required for parsing, or refused to produce a distribution. See Table 17.
A.3 Individual Sociodemographic Results
Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 8, 9, and 10 respectively.
Demographic Claude-3.5-v2 Llama-3.2-90B Mistral-large OLMo-2-7B-I Qwen-2.5-72B Average
$V$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $V_{C}$
Age 15-29 95.2 92.7 85.0 94.1 92.9 92.5 56.6 88.6 90.2 90.7 84.0 91.7
30-49 92.7 90.7 85.7 93.1 93.1 92.7 58.9 87.0 90.6 89.8 84.2 90.7
50+ 88.6 87.5 87.9 91.0 91.4 90.7 63.7 85.9 90.6 87.1 84.4 88.4
Edu. primary 83.9 83.8 78.0 86.3 84.6 84.0 62.1 88.6 82.5 82.6 78.2 85.1
secondary 95.1 91.2 86.5 91.8 92.1 92.2 57.0 85.8 91.1 90.5 84.4 90.3
tertiary 86.2 82.7 88.6 88.4 90.4 89.5 63.8 86.2 94.6 90.6 84.7 87.5
Employment full time for employer 92.3 88.5 89.1 91.1 93.4 92.7 60.6 86.0 93.9 90.8 85.9 89.8
full time for self 91.0 89.7 84.3 91.3 91.5 91.2 60.0 88.2 89.3 88.8 83.2 89.8
part time (no full time) 91.2 89.9 86.0 93.2 92.4 92.1 60.9 88.1 90.2 89.2 84.1 90.5
part time (full time) 92.1 90.6 83.1 93.0 90.3 90.4 58.5 89.7 88.0 87.9 82.4 90.3
out of work force 91.1 89.6 85.6 92.1 91.5 91.1 60.5 86.7 89.7 88.6 83.7 89.6
unemployed 92.9 92.7 80.3 91.4 88.1 88.8 51.6 87.4 85.0 87.5 79.6 89.6
Sex female 92.2 90.5 84.6 92.2 91.8 91.6 59.0 87.4 89.3 88.8 83.4 90.1
male 91.9 89.5 87.8 93.4 93.8 92.6 60.6 86.8 92.5 90.5 85.3 90.6
Income fourth 20% 92.9 90.7 87.5 93.3 94.3 93.2 60.1 86.9 92.0 90.3 85.4 90.9
middle 20% 91.9 90.3 86.2 92.5 92.7 92.4 60.2 87.1 90.7 89.4 84.3 90.3
poorest 20% 90.2 88.8 82.5 91.3 89.2 89.1 59.0 87.9 86.9 87.2 81.6 88.9
second 20% 91.0 89.5 85.0 91.7 91.4 91.1 60.4 87.3 89.4 88.4 83.4 89.6
top 20% 92.5 89.1 87.5 92.5 93.6 92.7 59.5 86.8 93.7 92.0 85.4 90.6
Area city/suburb 94.2 91.4 85.9 92.7 93.0 93.2 56.4 86.3 90.4 90.3 84.0 90.8
rural/small town 90.3 89.2 86.2 92.0 92.0 91.5 62.1 87.8 90.2 88.5 84.2 89.8
Table 8: Opinion alignment before ($V$) and after ($V_{C}$) calibration for WGM demographics using base-prompted, verbalized elicitation. Each pair of columns compares the base-generated distributions to the calibrated distributions (C), with significant differences between the two bolded. The two “Average” columns on the right are averages across models.
Demographic Claude-3.5-v2 Llama-3.2-90B Mistral-large OLMo-2-7B-I Qwen-2.5-72B Average
$V$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $V_{C}$
Age 18-29 92.5 92.6 87.5 88.8 86.9 88.0 68.9 83.2 89.7 89.2 85.1 88.4
30-49 92.6 92.5 86.5 87.8 85.9 87.1 66.6 81.5 89.2 88.7 84.2 87.5
50-64 92.5 92.4 86.5 87.8 84.1 85.8 63.5 78.6 88.1 87.5 82.9 86.4
65+ 92.1 92.1 87.3 88.3 83.9 85.4 62.1 76.9 87.9 87.3 82.7 86.0
Edu. Associate’s degree 92.5 92.4 85.9 87.4 84.6 86.5 65.3 80.5 88.3 87.6 83.3 86.9
College graduate 92.1 92.0 87.6 88.6 85.7 87.0 65.0 79.8 88.6 88.0 83.8 87.1
High school graduate 91.5 91.4 85.2 86.5 83.2 84.8 63.6 78.5 87.8 87.2 82.3 85.7
Less than high school 88.8 88.3 84.4 85.5 82.7 84.0 69.2 81.3 87.0 86.5 82.4 85.1
Postgraduate 92.2 92.1 87.3 88.4 85.4 86.3 64.0 78.7 88.2 87.6 83.4 86.6
Some college, no degree 93.7 93.7 87.0 88.4 84.9 86.5 65.2 80.2 89.4 88.8 84.0 87.5
Region Midwest 92.3 92.3 86.2 87.6 84.7 86.2 64.2 79.3 88.3 87.7 83.1 86.6
Northeast 92.9 92.8 88.2 89.1 85.6 87.1 64.0 78.7 89.1 88.6 84.0 87.3
South 92.9 92.8 87.1 88.3 84.9 86.6 65.1 79.9 88.9 88.4 83.8 87.2
West 92.8 92.7 87.9 89.1 86.0 87.3 65.0 80.1 89.1 88.5 84.2 87.5
Income \$100,000 or more 90.7 90.6 85.7 86.9 84.1 85.3 64.0 78.6 86.9 86.2 82.3 85.5
\$30,000-\$50,000 93.2 93.1 86.3 87.6 84.9 86.5 65.1 80.0 89.4 88.8 83.8 87.2
\$50,000-\$75,000 93.2 93.2 87.1 88.3 85.4 86.8 64.9 79.7 89.1 88.5 83.9 87.3
\$75,000-\$100,000 93.0 92.8 87.9 89.2 86.1 87.5 65.6 80.4 89.4 88.8 84.4 87.7
Less than \$30,000 93.9 93.7 88.4 89.8 86.1 87.7 65.7 80.7 90.2 89.5 84.9 88.3
Marital Divorced 94.3 94.1 86.8 88.3 85.0 86.8 63.7 78.8 89.5 88.7 83.9 87.3
Married 92.0 91.9 86.5 87.8 84.7 86.1 64.6 79.5 88.3 87.7 83.2 86.6
Never been married 93.3 93.2 88.5 89.9 87.5 88.5 66.5 81.4 89.9 89.3 85.1 88.5
Separated 93.8 93.6 89.2 90.3 86.4 87.6 65.4 79.4 90.1 89.7 85.0 88.1
Widowed 90.1 90.0 87.5 88.4 82.9 84.1 61.5 75.8 86.1 85.5 81.6 84.8
Pol. Party Democrat 90.5 91.0 89.5 89.2 85.3 85.8 62.5 77.3 87.9 88.4 83.1 86.3
Independent 92.1 92.0 86.3 87.4 84.3 85.7 64.6 79.6 88.2 87.6 83.1 86.5
Other 91.2 91.0 84.5 86.1 84.0 85.6 65.8 80.9 87.5 86.7 82.6 86.1
Republican 88.3 88.2 81.9 83.2 80.6 82.2 64.7 79.5 85.4 84.9 80.2 83.6
Race Asian 93.9 93.8 88.3 89.8 89.8 90.8 73.1 86.9 92.6 92.0 87.5 90.7
Black 93.4 93.6 91.4 91.8 87.4 88.1 65.7 80.3 90.5 90.0 85.7 88.8
Hispanic 94.3 94.3 90.3 91.2 87.6 88.6 67.1 81.9 91.6 91.1 86.2 89.4
Other 93.1 92.5 86.5 88.3 85.9 87.7 67.2 81.6 90.4 89.5 84.6 87.9
White 92.2 92.1 86.4 87.7 84.5 85.9 63.8 78.8 88.2 87.6 83.0 86.4
Religion Agnostic 89.5 89.5 85.1 86.4 82.9 84.1 63.3 77.4 85.7 84.9 81.3 84.5
Atheist 87.6 88.1 85.3 85.9 82.4 83.2 62.4 76.8 84.6 84.1 80.5 83.6
Buddhist 90.1 90.2 90.2 89.9 86.4 86.5 66.4 81.0 88.5 88.5 84.3 87.2
Hindu 88.3 88.6 86.3 85.9 86.9 86.4 73.0 85.2 87.2 87.5 84.3 86.7
Jewish 93.3 93.6 87.7 88.7 87.2 88.2 66.8 80.7 89.9 89.5 85.0 88.1
Mormon 87.5 87.2 82.6 84.0 82.9 84.2 67.8 81.4 85.8 85.2 81.3 84.4
Muslim 91.9 91.6 91.2 92.3 89.4 90.7 69.9 83.9 91.4 90.7 86.8 89.8
Nothing in particular 93.2 93.1 89.2 90.3 86.6 88.0 65.1 80.2 89.4 89.0 84.7 88.1
Orthodox 93.4 93.1 88.5 89.9 88.2 89.3 69.9 84.1 91.8 91.2 86.4 89.5
Other 92.7 92.5 86.1 87.6 84.0 85.6 63.4 78.5 89.5 88.9 83.1 86.6
Protestant 91.4 91.4 85.1 86.3 83.3 85.0 63.9 78.8 87.3 86.8 82.2 85.7
Roman Catholic 93.1 93.0 87.5 88.7 85.7 87.2 66.3 80.8 89.3 88.7 84.4 87.7
Sex Female 93.4 93.3 87.8 89.1 85.1 86.5 63.5 78.4 89.2 88.7 83.8 87.2
Male 91.5 91.4 86.1 87.6 85.4 86.6 66.2 81.1 88.1 87.4 83.5 86.8
Table 9: Opinion alignment before ($V$) and after ($V_{C}$) calibration for OQA demographics using base-prompted, verbalized elicitation. Each pair of columns compares the base-generated distributions to the calibrated distributions (C), with significant differences between the two bolded. The two “Average” columns on the right are averages across models.
Demographic Claude-3.5-v2 Llama-3.2-90B Mistral-large OLMo-2-7B-I Qwen-2.5-72B Average
$V$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $_{C} VV$ $V_{C}$
Age 16-24 years 76.4 81.6 65.5 81.6 73.9 81.6 86.2 88.3 75.3 79.2 75.5 82.5
25-34 years 76.8 81.8 65.9 81.8 74.3 81.8 86.6 88.8 75.6 79.4 75.8 82.7
35-44 years 76.7 81.7 65.9 81.7 74.1 81.7 87.1 88.2 75.2 79.2 75.8 82.5
45-54 years 76.0 81.0 65.5 81.0 73.6 81.0 87.5 87.6 74.6 78.5 75.4 81.8
55-64 years 74.6 79.5 64.3 79.5 72.3 79.5 87.7 85.8 73.2 77.1 74.4 80.3
65+ years 72.9 77.9 62.8 77.9 70.8 77.9 87.5 84.6 71.6 75.5 73.1 78.8
Education bachelor 74.1 79.1 63.7 79.1 71.8 79.1 86.9 85.3 72.8 76.7 73.9 79.9
doctoral 73.3 78.8 62.6 78.8 71.0 78.8 86.0 85.7 72.5 76.4 73.1 79.7
early childhood 78.7 83.7 68.1 83.7 76.4 83.7 90.2 91.3 77.4 81.2 78.2 84.7
lower secondary 75.6 80.6 65.3 80.6 73.4 80.6 87.6 87.1 74.3 78.2 75.2 81.4
master 74.3 79.5 63.6 79.5 72.0 79.5 86.8 87.0 73.2 77.1 74.0 80.5
post-secondary 75.4 80.4 64.6 80.4 72.8 80.4 86.2 86.2 74.0 78.0 74.6 81.1
primary 76.1 81.1 65.7 81.1 73.9 81.1 88.8 87.7 74.8 78.7 75.9 81.9
short-cycle tertiary 73.7 78.7 63.2 78.7 71.4 78.7 85.9 85.2 72.3 76.2 73.3 79.5
upper secondary 75.4 80.3 64.6 80.3 72.9 80.3 86.2 86.7 74.0 77.9 74.6 81.1
Employment full time 75.1 80.1 64.7 80.1 72.8 80.1 87.1 86.5 73.7 77.7 74.7 80.9
housewife 75.7 80.7 65.4 80.7 73.5 80.7 87.5 87.9 74.4 78.3 75.3 81.7
other 74.8 79.8 64.7 79.8 72.7 79.8 87.1 83.9 73.6 77.4 74.6 80.1
part time 76.8 81.8 66.2 81.8 74.3 81.8 87.1 88.2 75.3 79.4 75.9 82.6
retired/pensioned 72.9 77.9 62.6 77.9 70.7 77.9 87.1 84.8 71.6 75.4 73.0 78.8
self-employed 77.0 82.0 66.0 82.0 74.4 82.0 87.4 89.6 75.7 79.6 76.1 83.0
student 75.6 80.8 64.9 80.8 73.1 80.8 86.1 87.0 74.4 78.4 74.8 81.6
unemployed 78.0 83.0 67.0 83.0 75.4 83.0 86.9 89.0 76.8 80.6 76.8 83.7
Household size 1 74.1 79.1 63.8 79.1 72.0 79.1 87.2 85.0 72.9 76.7 74.0 79.8
2 73.5 78.5 63.3 78.5 71.4 78.5 87.0 84.6 72.2 76.1 73.5 79.2
3 75.0 80.0 64.5 80.0 72.6 80.0 87.1 85.9 73.4 77.5 74.5 80.7
4 75.6 80.5 65.1 80.5 73.1 80.5 86.9 86.7 74.0 78.1 74.9 81.3
5 76.5 81.4 65.5 81.4 73.9 81.4 87.0 88.8 75.2 79.0 75.6 82.4
6 76.8 81.8 65.7 81.8 74.1 81.8 87.2 89.5 75.6 79.4 75.9 82.9
7 persons or more 78.9 86.5 66.1 86.5 74.6 86.5 80.6 86.0 78.4 84.8 75.7 86.1
Imm. Immigrant 71.7 76.7 61.3 76.7 69.5 76.7 86.0 83.8 70.4 74.2 71.8 77.6
Native 76.1 81.1 65.5 81.1 73.5 81.1 87.0 87.6 74.6 78.7 75.3 81.9
Marital Divorced 74.2 79.2 63.9 79.2 72.0 79.2 86.7 84.8 72.9 76.8 73.9 79.8
Living together 74.1 79.2 63.5 79.2 71.7 79.2 83.8 82.7 72.6 76.7 73.1 79.4
Married 76.2 81.2 65.6 81.2 73.6 81.2 87.4 88.3 74.7 78.8 75.5 82.1
Separated 74.5 79.4 63.9 79.4 72.1 79.4 84.9 83.6 73.1 77.0 73.7 79.8
Single 75.9 81.0 65.3 81.0 73.5 81.0 86.6 87.1 74.5 78.6 75.2 81.7
Widowed 75.0 80.0 64.2 80.0 72.5 80.0 87.0 86.5 73.7 77.5 74.5 80.8
Religion Buddhist 74.8 79.8 63.8 79.8 72.1 79.8 85.8 86.1 73.5 77.3 74.0 80.6
Hindu 79.7 86.0 68.4 86.0 76.9 86.0 85.3 93.4 79.6 83.6 78.0 87.0
Jew 76.3 81.5 65.2 81.5 73.6 81.5 88.0 87.9 75.2 79.0 75.7 82.3
Muslim 76.9 81.8 66.1 81.8 74.4 81.8 89.5 90.5 75.6 79.4 76.5 83.1
Orthodox 73.1 78.6 62.3 78.6 70.7 78.6 85.5 85.8 72.4 76.2 72.8 79.6
Other 70.8 75.7 61.4 75.7 68.3 75.7 85.7 80.1 69.2 73.3 71.1 76.1
Other Christian 71.6 76.6 61.5 76.6 69.5 76.6 84.0 80.6 70.3 74.2 71.4 76.9
Protestant 74.3 79.3 63.7 79.3 71.9 79.3 85.3 83.9 72.9 76.9 73.6 79.7
Roman Catholic 73.9 78.9 63.4 78.9 71.6 78.9 85.2 83.8 72.6 76.5 73.3 79.4
none 70.6 75.6 61.0 75.6 68.5 75.6 84.9 81.1 69.3 73.2 70.9 76.2
Sex Female 75.2 80.2 64.3 80.2 72.5 80.2 86.1 86.5 73.8 77.7 74.4 81.0
Male 76.3 81.2 65.9 81.2 73.9 81.2 87.8 87.7 74.8 78.8 75.7 82.0
Table 10: Opinion alignment before ($V$) and after ($V_{C}$) calibration for WVS demographics using base-prompted, verbalized elicitation. Each pair of columns compares the base-generated distributions to the calibrated distributions (C), with significant differences between the two bolded. The two “Average” columns on the right are averages across models.
A.4 Out-of-domain generalization
We study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset. We train regression models on two of our three datasets and evaluate on the held-out dataset. Results are shown in Tables 11 and 12. As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset. Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM. This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
Model Base prompt Sociodemographic prompt
$P$ $P_{C}$ $S$ $S_{C}$ $V$ $V_{C}$ $P$ $P_{C}$ $S$ $S_{C}$ $V$ $V_{C}$
WGM OLMo-2-7B-Base 65.6 82.4 79.2 81.8 82.2 82.6 74.8 81.9 71.0 82.1 74.9 82.6
OLMo-2-7B-SFT 80.6 81.8 70.3 81.3 68.6 83.9 75.6 82.4 70.2 82.2 65.4 82.7
OLMo-2-7B-DPO 71.1 81.7 63.2 82.1 63.9 83.1 72.2 82.7 68.7 82.4 63.8 83.0
OLMo-2-7B-Instruct 59.6 82.0 67.7 81.8 59.8 82.8 62.3 81.7 64.5 82.2 69.8 84.4
Llama-3-70B 64.9 83.6 73.7 81.8 84.5 86.5 66.5 84.6 61.7 84.1 84.4 87.1
Llama-3.1-70B 73.2 82.7 70.3 84.4 80.6 81.8 68.5 85.2 66.1 81.8 85.2 87.9
Llama-3.2-1B 81.8 81.8 81.8 81.8 61.9 82.3 46.7 80.7 61.4 86.4
Llama-3.2-11B 73.7 85.9 73.2 81.4 73.6 82.1 71.8 84.5 72.0 84.1 71.7 81.1
Llama-3.2-90B 68.1 84.1 73.0 84.4 84.8 86.2 70.6 84.7 67.6 84.4 85.0 86.4
Qwen-2.5-72B 57.7 83.5 53.3 83.0 88.2 86.7 66.0 81.8 63.4 83.7 89.1 85.9
Mistral-small 61.6 84.1 61.6 83.7 88.9 85.7 67.5 84.6 64.4 83.9 89.0 85.0
Mistral-large 62.4 84.4 72.0 83.5 89.4 86.8 68.4 84.6 63.0 84.2 87.3 86.2
Claude-3 57.8 82.2 50.0 81.8 84.8 84.3 64.9 82.2 52.9 81.6 88.4 85.6
Claude-3.5-v1 64.5 83.9 62.0 83.5 86.1 84.9 65.8 84.7 62.3 84.5 86.3 88.1
Claude-3.5-v2 66.2 84.3 59.5 83.1 89.3 86.5 65.4 84.3 61.4 84.1 89.0 87.7
Average 67.3 83.2 67.4 82.6 80.3 84.6 68.1 83.5 63.7 83.1 79.4 85.3
Std Dev 7.5 1.2 9.0 1.1 9.9 1.8 4.1 1.3 6.7 1.2 10.3 2.2
OQA OLMo-2-7B-Base 85.0 89.1 82.6 88.4 82.1 88.0 80.8 88.3 77.2 88.3 83.8 88.4
OLMo-2-7B-SFT 78.8 88.4 79.4 88.4 68.7 83.4 80.7 88.3 77.8 88.4 72.6 84.2
OLMo-2-7B-DPO 79.3 88.4 81.5 88.3 69.3 83.6 82.1 88.2 78.7 88.3 72.3 85.4
OLMo-2-7B-Instruct 72.6 88.6 72.3 88.4 65.4 82.5 72.8 88.5 70.5 88.6 68.5 81.4
Llama-3-70B 72.2 88.4 76.8 88.4 81.8 88.4 74.4 88.7 70.1 89.0 79.5 88.8
Llama-3.1-70B 73.5 88.9 70.5 88.9 83.8 88.4 76.4 89.1 71.7 90.4 83.6 89.6
Llama-3.2-1B 82.3 88.4 88.4 88.4 72.6 89.1 70.2 88.5 83.6 87.6
Llama-3.2-11B 84.3 88.4 75.5 87.7 65.8 88.2 75.7 87.8 74.8 87.7 83.0 88.7
Llama-3.2-90B 79.3 89.4 75.3 88.4 86.8 88.4 79.2 88.9 76.1 89.0 83.4 88.8
Qwen-2.5-72B 73.9 88.4 67.0 88.5 88.4 88.9 74.4 88.4 71.3 88.4 89.2 89.1
Mistral-small 77.3 88.4 73.6 88.4 86.9 88.4 75.1 88.4 71.6 88.4 87.8 89.5
Mistral-large 79.5 88.8 75.3 88.4 85.0 88.8 75.8 88.9 72.4 88.4 83.8 87.5
Claude-3 73.5 88.4 66.4 88.4 87.6 88.9 76.4 88.4 72.6 88.5 89.3 89.3
Claude-3.5-v1 70.5 88.4 69.6 88.4 89.4 88.4 76.1 89.5 72.8 89.7 91.0 91.5
Claude-3.5-v2 70.5 88.4 72.5 88.4 91.7 89.2 76.1 89.2 73.3 89.4 91.9 90.8
Average 76.8 88.6 75.1 88.4 80.9 87.4 76.6 88.6 73.4 88.7 82.9 88.0
Std Dev 4.8 0.3 6.0 0.2 9.4 2.3 2.9 0.5 2.8 0.7 7.0 2.6
WVS OLMo-2-7B-Base 74.5 81.2 70.7 80.3 84.5 84.4 78.2 81.1 78.2 81.0 80.8 83.0
OLMo-2-7B-SFT 69.7 79.1 68.1 79.2 87.2 85.6 72.9 80.3 73.1 80.8 84.9 87.0
OLMo-2-7B-DPO 66.2 78.3 70.5 79.7 89.1 82.8 70.6 80.3 70.2 80.8 82.5 84.3
OLMo-2-7B-Instruct 75.3 80.9 74.9 80.5 86.6 87.6 58.3 78.1 60.0 76.9 86.0 88.6
Llama-3-70B 59.7 76.3 61.7 79.3 81.3 82.9 67.9 79.2 62.8 79.6 78.1 79.4
Llama-3.1-70B 61.9 76.8 56.4 75.8 65.2 70.7 63.6 78.0 57.6 76.7 68.3 72.1
Llama-3.2-1B 80.3 80.3 80.3 80.3 68.7 82.5 66.6 79.0
Llama-3.2-11B 75.4 81.9 76.0 84.4 82.7 83.6 78.2 83.2 78.0 81.9 66.5 77.0
Llama-3.2-90B 61.6 76.8 59.1 75.4 64.6 70.2 62.1 77.7 59.5 76.5 67.7 71.5
Qwen-2.5-72B 39.6 71.7 42.4 74.3 74.0 74.3 49.1 71.8 49.4 74.7 73.4 73.8
Mistral-small 42.9 73.0 46.3 74.4 75.0 75.8 48.0 72.5 46.4 74.6 68.6 71.0
Mistral-large 48.8 74.2 44.0 76.4 72.8 74.8 54.3 74.4 51.5 76.4 76.6 78.3
Claude-3 47.3 71.4 55.7 76.3 74.1 74.0 55.3 74.8 57.3 77.1 73.2 73.3
Claude-3.5-v1 44.3 71.2 49.6 74.6 75.7 77.7 58.7 76.7 54.8 78.1 73.3 75.5
Claude-3.5-v2 46.5 72.1 51.0 75.5 75.2 75.2 61.0 78.2 56.8 78.4 75.6 75.6
Average 59.6 76.3 60.4 77.8 77.7 78.5 63.1 77.9 61.5 78.2 75.4 77.9
Std Dev 13.8 3.9 12.4 3.0 7.7 5.8 9.5 3.4 9.9 2.3 6.4 5.8
Table 11: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method, when training on two datasets and evaluating on the out-of-domain dataset. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase ($P$), self-random ($S$), and verbalized ($V$). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
Model Base prompt SD prompt
$L$ $_{C} LL$ $L_{C}$
WGM Llama-3-70B 73.1 84.6 67.5 86.1
Llama-3.1-70B 74.0 81.8 70.4 86.1
Llama-3.2-1B 83.6 83.5 83.0 83.1
Llama-3.2-11B 84.0 84.3 82.2 81.8
Average 78.7 83.5 75.8 84.3
Std Dev 5.9 1.3 8.0 2.2
OQA Llama-3-70B 75.4 88.4 72.3 88.5
Llama-3.1-70B 78.9 88.4 75.7 88.4
Llama-3.2-1B 88.3 85.2 87.3 86.6
Llama-3.2-11B 82.6 88.3 88.3 88.4
Average 81.3 87.6 80.9 88.0
Std Dev 5.5 1.6 8.1 0.9
WVS Llama-3-70B 54.1 74.8 51.1 72.1
Llama-3.1-70B 43.5 71.3 49.1 70.8
Llama-3.2-1B 83.9 82.1 85.1 84.6
Llama-3.2-11B 74.8 78.3 72.9 75.9
Average 64.1 76.6 64.5 75.8
Std Dev 18.5 4.6 17.4 6.2
Table 12: Opinion alignment before and after calibration for each dataset and LLM using log probability distributions, when training on two datasets and evaluating on the out-of-domain dataset. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for log probabilities ($L$). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
Appendix B Prompts (Reproducibility)
B.1 Demographics used
Demographics are shown in Table 13.
Dataset Qs Example question Demographics
WGM 14 How much do you, personally, know about science? Do you know a lot, some, not much, or nothing at all?
How much do you trust doctors and nurses in your country?
age (15-29, 30-49, 50+), education(primary, secondary, tertiary), employment status (employed full time for an employer, employed full time for self, employed part time but do not want full time, unemployed, employed part time want full time, out of work force), income quintile (poorest 20%, second 20%, middle 20%, fourth 20%, top 20%), living area (city, rural), sex (male, female), world region (Eastern Africa, Central Africa, North Africa, Southern Africa, Western Africa, Central America and Mexico, Northern America, South America, Central Asia, East Asia, Southeast Asia, South Asia, Middle East, Eastern Europe, Northern Europe, Southern Europe, Western Europe, Aus/NZ)
OQA 38 How important, if at all, is being a gun owner to your overall identity?
How much confidence, if any, do you have in elected officials to act in the best interests of the public?
age (18-29, 30-49, 50-64, 65+), education (Less than high school, High school graduate, Some college, no degree, Associates degree, College graduate/some postgrad, Postgraduate), income (Less than \$30,000, \$30,000-\$50,000, \$50,000-\$75,000, \$75,000-\$100,000, \$100,000 or more), marital status (Married, Divorced, Separated, Widowed, Never been married), political ideology (Very conservative, Conservative, Moderate, Liberal, Very liberal), political party (Republican, Democrat, Independent, Other), race (White, Black, Asian, Hispanic, Other), US region (Northeast, Midwest, South, West), religion (Protestant, Roman Catholic, Mormon, Orthodox, Jewish, Muslim, Buddhist, Hindu, Atheist, Agnostic, Other, Nothing in particular), sex (Male, Female)
WVS 40 Do you agree strongly, agree, disagree, or disagree strongly with the following statement? One of my main goals in life has been to make my parents proud.
Do you agree, disagree or neither agree nor disagree with the following statement? When jobs are scarce, employers should give priority to people of this country over immigrants.
age (16-24 years, 25-34 years, 35-44 years, 45-54 years, 55-64 years, 65+ years), education (early childhood, primary, lower secondary, upper secondary, post-secondary, short-cycle tertiary, bachelor or equivalent, master or equivalent, doctoral or equivalent), employment status (full time, part time, self-employed, retired/pensioned, housewife not otherwise employed, student, unemployed), household size (1, 2, 3, 4, 5, 6, 7 persons or more), immigrant status (I am born in this country, I am an immigrant to this country), income (low, middle, high), marital status (Married, Living together as married, Divorced, Separated, Widowed, Single), religion (none, Roman Catholic, Protestant, Orthodox (Russian/Greek/etc.), Jew, Muslim, Hindu, Buddhist, Other Christian), sex (Male, Female)
Table 13: The number of questions (Qs) and two example questions from each dataset, along with the demographics we use for each dataset. We do not use all the questions or demographics available due to cost for generations for each prompt, model, and elicitation method (270 settings, see Section 3.1).
B.2 Base prompts
Base prompts are shown in Table 14.
Prompt type Prompt template
Verbalized What percentage of people would respond with {answer choices} to the following question? Question: ‘{q}’ Respond with a list of {l} numbers which represent the percentages for {answer choices} in that order. Expected response format: [, , , ,] Response:
Log probabilities Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: {answer}
Self-random Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 1 Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 2 How might someone respond to the following question? Question: ‘{q}’ Possible responses are {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 3 Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 4 Below you will be asked to answer some questions. Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 5 Answer the following question. Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Table 14: Prompt templates used for base prompts (without any sociodemographic information). {q} is the question from the original dataset with minor modifications to be suitable for LLMs. {answer chocies} are the likert-scale answers to the questions from the original dataset.
B.3 Sociodemographic prompts
Sociodemographic prompts are shown in Table 15.
Prompt type Prompt template
Verbalized What percentage of people whose {att} is {c} would respond with {answer choices} to the following question? Question: ’{q}’ Respond with a list of {l} numbers which represent the percentages for {answer choices} in that order. Expected response format: [, , , ,] Response:
Log probabilities Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: {answer}
Self-random Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 1 Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 2 How would a person whose {att} is {c} respond to the following question? Question: ‘{q}’ Possible responses are {answer choices } Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 3 Person: {c} {att} Question: ‘{q}’ Response options: {answer choices } Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 4 Below you will be asked to provide a short description of your {att} and then answer some questions. Description: I am {c} Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Paraphrase 5 Answer the following question as if your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response:
Table 15: Prompt templates used for sociodemographic prompts (with sociodemographic information). {q} is the question from the original dataset with minor modifications to be suitable for LLMs. {answer chocies} are the likert-scale answers to the questions from the original dataset. {att} and {c} correspond to the demographic attribute and class respectively (e.g., “age” and “15-24 years”). We note that the placement of {att} and {c} in the prompt might be slightly different/inverted depending on the demographic for correct grammar.
B.4 Prompt variations
Prompt variations are shown in Table 16.
Prompt variation Prompt template
Few shot Please complete the task below. Follow the examples given. Examples: {5 examples} Task: {Original prompt}
CoT {Original question} Explain your reasoning step-by-step before answering. {Response instructions} Put your final answer at the end of your response between <ANS_START> and </ANS_END>. Use a maximum of 500 words for your reasoning and final answer combined.
Table 16: Prompt templates for the different prompt variants. This text is combined with the templates for both base prompts (Table 14) and SD prompts (Table 15) to get the results shown in Tables 2 and 4. All other results use the standard prompts without variation.
Appendix C Extracted distributions
The percent of distributions extracted from each LLM for each dataset and elicitation method is shown in Table 17.
Model Base prompt SD prompt
$P$ $S$ $V$ $P$ $S$ $V$
WGM Claude-3 100 100 100 100 100 100
Claude-3.5-v1 100 100 100 100 100 100
Claude-3.5-v2 100 100 100 100 100 100
Llama-3-70B 100 100 100 100 100 100
Llama-3.1-70B 100 100 82.4 100 100 92.8
Llama-3.2-1B 100 100 0 100 100 1.3
Llama-3.2-11B 100 100 70.6 100 100 27.1
Llama-3.2-90B 100 100 100 100 100 94.9
Mistral-small 100 100 100 100 100 99.7
Mistral-large 100 100 100 100 100 100
OLMo-2-7B-B 100 100 47.1 100 100 69.3
OLMo-2-7B-S 100 100 100 100 100 97.6
OLMo-2-7B-D 100 100 100 100 100 92.9
OLMo-2-7B-I 100 100 70.6 100 100 89.1
Qwen-2.5-72B 100 100 100 100 100 100
OQA Claude-3 100 100 100 100 100 100
Claude-3.5-v1 100 100 100 100 100 100
Claude-3.5-v2 100 100 100 100 100 100
Llama-3-70B 100 100 91.9 100 100 99.8
Llama-3.1-70B 100 100 100 100 12.1 93.4
Llama-3.2-1B 100 100 0 100 100 0.6
Llama-3.2-11B 100 100 32.4 100 100 7.2
Llama-3.2-90B 100 100 100 100 100 99.7
Mistral-small 100 100 100 100 100 100
Mistral-large 100 100 100 100 100 99.7
OLMo-2-7B-B 100 100 100 100 100 99.9
OLMo-2-7B-S 100 100 100 100 100 99.5
OLMo-2-7B-D 100 100 94.6 100 100 79.8
OLMo-2-7B-I 100 100 75.7 100 100 79.5
Qwen-2.5-72B 100 100 100 100 100 99.4
WVS Claude-3 100 100 100 100 100 100
Claude-3.5-v1 100 100 100 19.6 19.6 19.6
Claude-3.5-v2 100 100 100 49 49 49
Llama-3-70B 100 100 97.9 99.4 99 87.2
Llama-3.1-70B 100 100 100 49 49 47.2
Llama-3.2-1B 100 100 0 49 49 0.4
Llama-3.2-11B 100 100 39.7 49 49 3.7
Llama-3.2-90B 100 100 100 100 100 98.3
Mistral-small 100 100 100 100 100 100
Mistral-large 100 100 100 100 100 99.9
OLMo-2-7B-B 100 100 100 100 100 99.6
OLMo-2-7B-S 100 100 100 100 100 99.3
OLMo-2-7B-D 100 100 76.8 100 100 72.2
OLMo-2-7B-I 100 100 82.1 100 100 79.3
Qwen-2.5-72B 100 100 100 100 100 98.6
Table 17: The percent of extracted distributions in each setting. We found that smaller models (especially Llama 3.2 1B and 11B) would often refuse to answer questions (question would hit safety restrictions for role-playing as a particular identity group), or did not follow instructions in formatting output distributions.